### v1

# NAP-Tuning: Neural Augmented Prompt Tuning for Adversarially Robust Vision-Language Models

Abstract
Vision-Language Models (VLMs) such as CLIP have demonstrated remarkable capabilities in understanding relationships between visual and textual data through joint embedding spaces. Despite their effectiveness, these models remain vulnerable to adversarial attacks, particularly in the image modality, posing significant security concerns. Building upon our previous work on Adversarial Prompt Tuning (AdvPT), which introduced learnable text prompts to enhance adversarial robustness in VLMs without extensive parameter training, we present a significant extension by introducing the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). Our key innovations include: (1) extending AdvPT from text-only to multi-modal prompting across both text and visual modalities, (2) expanding from single-layer to multi-layer prompt architectures, and (3) proposing a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification to directly address the distortions introduced by adversarial attacks in feature space. Our NAP-Tuning approach incorporates token refiners that learn to reconstruct purified features through residual connections, allowing for modality-specific and layer-specific feature correction. Comprehensive experiments demonstrate that NAP-Tuning significantly outperforms existing methods across various datasets and attack types. Notably, our approach shows significant improvements over the strongest baselines under the challenging AutoAttack benchmark, outperforming them by 33.5% on ViT-B16 and 33.0% on ViT-B32 architectures while maintaining competitive clean accuracy. This work establishes the importance of architecture redesign in prompt tuning for adversarial robustness, moving beyond loss-focused approaches to create an adaptive defense mechanism that can identify and rectify adversarial perturbations across embedding spaces.
Index Terms:
Adversarial robustness, Vision-Language models, Prompt tuning, Feature purification, Neural augmentation
I Introduction
Large-scale pre-trained Vision-Language Models (VLMs) have demonstrated remarkable capabilities in understanding and connecting visual and textual information, enabling impressive performance on a wide range of downstream tasks. Models like CLIP [1] and ALIGN [2] have shown unprecedented zero-shot transfer abilities by learning from web-scale image-text pairs. As these models gain widespread adoption in real-world applications, ensuring their robustness against adversarial attacks becomes increasingly critical.
In our previous work [3], we introduced Adversarial Prompt Tuning (AdvPT), a novel approach that enhances the adversarial robustness of VLMs by aligning text embeddings with adversarial image embeddings through learnable text prompts. This method represented a paradigm shift from traditional adversarial training [4] approaches by focusing on prompt-level modifications rather than model parameter retraining, offering significant computational advantages while maintaining effectiveness.
While AdvPT demonstrated the potential of prompt tuning for adversarial defense, it exhibited three critical limitations: (1) its restriction to text modality prompts, (2) its reliance on single-layer prompting, and (3) its emphasis on loss function redesign rather than architectural innovations. Following our initial work, several approaches [5, 6] have attempted to address some of these issues, particularly by extending prompting to multiple modalities and layers. However, these incremental improvements fail to address the core architectural challenge of prompt tuning in adversarial settings.
The fundamental limitation across all existing approaches—including both our original AdvPT and subsequent works—is the direct transplantation of prompt tuning techniques from the generalization domain (e.g., CoOp [7] and MaPLe [8]) to adversarial defense without rethinking the underlying architecture. This approach overlooks a critical insight: prompt tuning, originally developed for clean data environments, faces fundamentally different challenges in adversarial scenarios where feature spaces are deliberately corrupted. Adversarial perturbations introduce systematic distortions that cannot be effectively countered through loss modifications alone, regardless of whether the prompting occurs in single or multiple modalities, or at one or multiple layers.
To comprehensively address all three limitations, we propose the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning). Our approach not only extends prompting to multiple modalities and layers but, most critically, fundamentally reconceptualizes prompt architecture for adversarial defense by integrating feature purification mechanisms. The Neural Augmentor framework incorporates token refiners—lightweight neural networks that learn to reconstruct clean feature representations from adversarially perturbed inputs through residual connections. This architecture enables modality-specific and layer-specific feature correction, yielding substantially enhanced adversarial robustness. Fig. LABEL:fig:comparison illustrates the evolution from text-only prompting to our comprehensive NAP-Tuning, demonstrating how our method facilitates clean feature reconstruction from adversarially corrupted inputs across both modalities and multiple network layers.
Through extensive experiments under rigorous evaluation protocols, we demonstrate that our approach significantly outperforms our initial AdvPT and other state-of-the-art methods, achieving superior robustness while maintaining competitive clean accuracy. Our analysis revealed that NAP-Tuning effectively learns to correct adversarial distortions in feature space, providing insights into the working mechanisms of successful adversarial defenses.
As an extension of our conference paper [3], this work makes several significant contributions:
  • Multi-modal defense framework: We extend the initial text-only AdvPT to a comprehensive multi-modal framework that incorporates learnable prompts in both text and visual pathways, creating a more effective defense mechanism that addresses vulnerabilities across modalities.
  • Multi-layer prompt architecture: We expand from single-layer prompting to multi-layer prompt architectures, allowing for more nuanced and layer-specific defenses that can address adversarial perturbations manifesting differently at various depths of the network.
  • Neural Augmentor design: Most significantly, we propose a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification via token refiners. This represents a paradigm shift from previous loss-focused prompt tuning methods to architecture-focused prompt tuning defense, addressing the fundamental limitations of existing approaches.
  • Comprehensive evaluation: We provide extensive experimental validation across multiple datasets and attack types under stricter evaluation settings. Our approach demonstrates significant performance improvements over the previous conference version (AdvPT), achieving average gains of 42.6% under PGD-100 attacks and 47.5% under the more challenging AutoAttack benchmark, establishing new standards for adversarial robustness in VLMs.
II Related Work
II-A Vision-Language Models
VLMs have established themselves as powerful frameworks for bridging visual and textual modalities. These models can be categorized into two principal architectures. The first category comprises generative models built upon Large Language Models (LLMs), such as LLaVA [9] and Qwen-VL [10], which extend text-only LLMs with vision encoders to enable multimodal understanding and generation.
The second category—which is the focus of adversarial prompt tuning techniques—utilizes contrastive learning to establish joint embeddings of images and text. Models such as CLIP [1] and ALIGN [2] project both modalities into a unified semantic space where related content exhibits higher similarity. These models demonstrate exceptional capabilities across diverse tasks through their dual-encoder architecture. However, their susceptibility to adversarial attacks, particularly in the visual domain, poses significant security concerns as their deployment expands into critical applications.
II-B Prompt Learning
Prompt learning emerged in natural language processing as an efficient parameter-efficient adaptation technique for pre-trained language models [11, 12]. This paradigm focuses on optimizing input transformations rather than modifying model parameters.
For VLMs, text-based prompt learning methods such as CoOp [7] and CoCoOp [13] demonstrated that replacing hand-crafted templates with learnable context vectors could significantly improve model performance. Our previous work, AdvPT [3], applied this concept to adversarial defense, showing that learnable text prompts could enhance robustness against adversarial attacks without modifying the underlying model. Multi-modal prompt learning approaches including MaPLe [8] and PromptKD [14] subsequently extended this paradigm by incorporating prompts in both textual and visual pathways. These architectural innovations have provided foundational techniques for multi-modal adversarial prompt tuning, though their primary focus on generalization rather than robustness presents limitations when applied to adversarial defense.
II-C Adversarial Prompt Tuning
Conventional adversarial defense methods such as adversarial training [4], TRADES [15], and MART [16] have proven effective but remain computationally prohibitive for large-scale VLMs. These approaches typically require training the whole model parameters from scratch, making them impractical for deployment in VLMs. An innovative alternative approach was introduced by Mao et al. [17] through their TeCoA framework, which intelligently leverages pre-trained VLM weights as initialization and selectively finetunes CLIP’s vision encoder.
Our previous work, AdvPT [3], introduced a more efficient and more effective paradigm by applying prompt learning to adversarial defense in VLMs. By aligning text embeddings with adversarial image embeddings through learnable prompts, AdvPT demonstrated that robust representations could be achieved with minimal computational overhead. Recent extensions to this approach include FAP [5], which leverages limited data to learn adversarially correlated text supervision while enhancing multi-modal feature consistency and uni-modal feature differentiation between natural and adversarial examples.
Recent advances in test-time optimization have yielded particularly impressive results in defending VLMs against adversarial attacks. Sheng et al. [18] developed R-TPT, an elegant test-time prompt tuning method that enhances the robustness of vision-language models by minimizing pointwise entropy and introducing a reliability-based weighted ensembling strategy. Their approach demonstrates the power of adaptive defenses that operate at inference time. TAPT [19] introduces an unsupervised approach that optimizes bimodal prompts at test time to enhance CLIP’s robustness while maintaining clean performance. In parallel, Xing et al. [20] introduced TTC, an innovative training-free defense mechanism that ingeniously leverages CLIP’s pre-trained vision encoder to counteract adversarial images during inference. Their approach offers complementary protection that can be effectively combined with existing robust models, demonstrating the value of ensemble strategies in comprehensive defense frameworks.
While these methods have achieved incremental improvements through loss function innovations, they overlook a critical consideration: the fundamental mismatch between prompt architectures designed for generalization and the unique requirements of adversarial robustness. Addressing adversarial vulnerabilities requires mechanisms specifically engineered to counteract feature-level distortions—a challenge that far exceeds general-purpose representation enhancement. This insight motivates our current work, which demonstrates that architectural innovations in prompt design can substantially improve adversarial robustness beyond what is possible through loss function modifications alone. By reimagining prompt architecture specifically for adversarial defense, we address the feature distortion mechanisms that underlie successful attacks on VLMs.
III Preliminary: Adversarial Prompt Tuning
To provide context for our extended approach, we first revisit the key concepts of Adversarial Prompt Tuning (AdvPT) as introduced in our previous work [3].
III-A Revisiting CLIP
We first provide a brief overview of VLMs, with a particular emphasis on the CLIP [1] architecture. Although our methods are specifically designed for CLIP, they are potentially extendable to a broader range of VLMs that are based on the contrastive learning paradigm. CLIP consists of an image encoder $f_{I}$ and a text encoder $f_{T}$, which project images and text into a shared embedding space. The model is trained to maximize the similarity between matched image-text pairs while minimizing the similarity between unmatched pairs.
For an image $x$ and a text description $t$, CLIP computes embeddings $f_{I}(x)$ and $f_{T}(t)$, and the similarity is calculated as:
$s(x,t)=\frac{f_{I}(x)\cdot f_{T}(t)}{||f_{I}(x)||\cdot||f_{T}(t)||}$ (1)
For classification tasks, CLIP computes the similarity between an image and a set of text templates describing each class (e.g., “a photo of a [CLASS]”), and selects the class with the highest similarity.
III-B Adversarial Attacks on VLMs
Adversarial attacks on VLMs can be categorized into two types based on the attacker’s access to model components. In our previous work [3], we considered a restricted threat model where the attacker only has access to the image encoder. Under this assumption, adversarial examples are crafted by maximizing the KL divergence between clean and perturbed image features:
$x_{adv}$ $=\arg\max_{x^{\prime}}D_{KL}(f_{I}(x),f_{I}(x^{\prime}))$
[…]
where $D_{KL}$ is the KL divergence measure. In this work, we consider a more challenging threat model that assumes the attacker has access to both image and text encoders, including textual prompt. This stronger adversary can generate more effective attacks by maximizing the cross-entropy loss between image-text pairs (even when the text is represented as learnable vectors):
$x_{adv}$ $=\arg\max_{x^{\prime}}\mathcal{L}_{CE}(f_{I}(x^{\prime}),f_{T}(t_{y}))$
[…]
where $t_{y}$ is the text template for class $y$, and $\mathcal{L}_{CE}$ is the cross-entropy loss measuring image-text similarity.
III-C Basic Adversarial Prompt Tuning
AdvPT addresses the vulnerability of VLMs to adversarial attacks by learning to align text embeddings with adversarial image embeddings through prompt tuning. The key insight is that this alignment can enhance robustness without requiring modifications to the model architecture. In AdvPT, the standard text template “a photo of a [CLASS]” is replaced with a template containing learnable context vectors:
[…]
where $\mathcal{V}_{t}$ are learnable vectors that are optimized to align with adversarial image embeddings, and $y$ represents class label. The training objective for AdvPT is to maximize the similarity between the clean text embedding (with learned prompt) and the adversarial image embedding:
$\min_{\mathcal{V}_{t}}\mathcal{L}(x_{adv},y)=\max_{\mathcal{V}_{t}}s(x_{adv},(\mathcal{V}_{t},y)),$ (5)
where $x_{adv}$ is the adversarial version of image $x$, and $(\mathcal{V}_{t},y)$ is the text template for the true class $y$.
IV Neural Augmentor for Multi-Modal Adversarial Prompt Tuning
Adversarial attacks fundamentally operate by introducing carefully crafted perturbations to input data that cause neural networks to make incorrect predictions. In the context of VLMs, these perturbations manipulate feature representations in ways that disrupt the crucial alignment between visual and textual information. Building upon our preliminary work on AdvPT, we introduce a comprehensive redesign of prompt-based adversarial defenses through our NAP-Tuning framework. This section details our approach, its architectural innovations, and theoretical foundations.
IV-A Framework Overview
Figure 3: Overview of our proposed Neural Augmentor module for multi-modal AdvPT.
Fig. LABEL:fig:comparison (c) introduces the overall framework, while Fig. 3 illustrates the architecture of our Neural Augmentor module. At its core, our approach aims to defend VLMs against adversarial attacks by addressing their fundamental vulnerability: the distortion of feature representations. Unlike previous prompt defenses that focus on loss function modifications or robust classification boundaries, our method directly targets the feature corruption mechanism through specialized architectural components.
The framework consists of three primary components that work in concert. First, a multi-modal prompting system introduces learnable prompt vectors that operate simultaneously in both textual and visual pathways of the model. Second, a multi-layer prompt architecture implements hierarchical prompting structures at different depths within the transformer architecture to address adversarial effects at various levels of feature abstraction. Third, Neural Augmentor modules serve as specialized neural networks that perform feature purification through token refiners.
IV-B Structural Innovations
IV-B1 Multi-Modal Prompting
A key limitation of our original AdvPT approach was its exclusive focus on the text modality, which left the visual pathway vulnerable to direct attacks. Our enhanced framework addresses this limitation by implementing learnable prompts in both textual and visual pathways, creating a coordinated defense system.
For the text pathway, we build upon the original AdvPT framework with enhanced prompt vectors $(\mathcal{V}_{t},y)$. For the visual pathway, we introduce a parallel set of prompt vectors that operate on the visual inputs $(\mathcal{V}_{i},x_{adv})$, where $\mathcal{V}_{i}$ are learnable visual prompt vectors. These visual prompts serve a complementary function to the text prompts, helping to guide perturbed visual representations back toward their clean manifold. Equation (5) can therefore be reformulated as:
$\min_{\mathcal{V}_{i},\mathcal{V}_{t}}\mathcal{L}(x_{adv},y)=\max_{\mathcal{V}_{i},\mathcal{V}_{t}}\ s((\mathcal{V}_{i},x_{adv}),(\mathcal{V}_{t},y)).$ (6)
The dual-modal prompt system enables a more comprehensive defense, as it can address attacks that target either modality or the cross-modal alignment. This is particularly important in VLMs, where adversarial perturbations can disrupt the crucial cross-modal matching that underlies the model’s performance.
IV-B2 Multi-Layer Prompt Architecture
Adversarial perturbations manifest differently across network depths, affecting both low-level perceptual features and high-level semantic representations. Our multi-layer prompt architecture addresses this by placing learnable prompts at multiple depths within the transformer network:
$h^{(l)}=\text{Layer}^{(l)}([\mathcal{V}^{(l)},h^{(l-1)}]),$ (7)
where $h^{l}$ represents the output of layer $l$, $\mathcal{V}^{(l)}$ are layer-specific learnable prompt vectors, and $\text{Layer}^{(l)}$ is the transformer layer function. This hierarchical prompting structure allows for targeted interventions at different levels of feature abstraction. Let $\mathbf{V}_{i}=\{\mathcal{V}_{i}^{j}\}_{j=1}^{l}$ and $\mathbf{V}_{t}=\{\mathcal{V}_{t}^{j}\}_{j=1}^{l}$, then Equation (6) can be reformulated as:
$\min_{\mathbf{V}_{i},\mathbf{V}_{t}}\mathcal{L}(x_{adv},y)=\max_{\mathbf{V}_{i},\mathbf{V}_{t}}s((\mathbf{V}_{i},x_{adv}),(\mathbf{V}_{t},y)).$ (8)
The multi-layer design offers several advantages over the single-layer approach of our original AdvPT. It enables depth-specific defense, where different layers can address distinct manifestations of adversarial perturbations, from low-level texture disruptions to high-level semantic shifts. It also facilitates progressive feature refinement, where corrections applied at earlier layers can be further refined by later layers, creating a cascading purification effect. More importantly, the increased parameter capacity allows the model to learn more complex defense strategies without modifying the underlying VLM weights.
This hierarchical structure represents a significant advancement over our preliminary work, which was limited to modifying only the textual prompts. By extending the defense mechanism throughout the network, we create a more robust barrier against adversarial perturbations that might otherwise bypass single-layer defenses.
IV-C Neural Augmentor Design
The most significant innovation in our framework is the Neural Augmentor—a specialized neural module designed explicitly for feature purification. Unlike conventional prompt tuning that focuses on optimizing context vectors, our Neural Augmentor actively transforms feature representations to counteract adversarial distortions.
IV-C1 Feature Distortion Theory
To formalize our approach, we first characterize the effect of adversarial perturbations in feature space. Let $\mathcal{X}$ denote the input space and $\mathcal{F}$ the feature space. For a clean image $x\in\mathcal{X}$ and its adversarial counterpart $x_{adv}=x+\delta$ , the feature distortion can be expressed as:
$\Delta f_{I}=f_{I}(x_{adv})-f_{I}(x),$ (9)
where $f_{I}(x)$ represents the feature representation of input $x$. Traditional adversarial defenses aim to make classification boundaries robust to these distortions. In contrast, our Neural Augmentor directly targets the distortion itself, attempting to recover:
$\hat{f}_{I}(x)\approx f_{I}(x_{adv})+\Phi(f_{I}(x_{adv})),$ (10)
where $\Phi$ is a learned correction function that approximates $-\Delta f$, effectively “purifying” the adversarial features (including both intermediate and final feature representations) back toward their clean counterparts.
IV-C2 TokenRefiner Architecture
The core component of the Neural Augmentor is the TokenRefiner $R$—a lightweight neural network that processes individual token representations to identify and correct adversarial distortions. For each token representation $z\in\mathbb{R}^{d}$, the TokenRefiner computes a corrective term $\tilde{z}=R(z)$. The TokenRefiner function is implemented as a two-layer network with a residual connection. It ensures stable gradient flow during training, facilitating effective optimization. Moreover, when no correction is needed (e.g., for clean inputs), the network can learn to output values near zero, effectively preserving the original features through an identity fallback mechanism. The $R$ operates on both text and visual tokens, with modality-specific parameters that allow it to learn distinct correction patterns for each pathway.
We apply the TokenRefiner $R$ to correct potentially perturbed feature representations before combining them with learnable vectors for the attention mechanism. Therefore, Equation (7) can be reformulated as:
$h^{(l)}=\text{Layer}^{(l)}\big{(}\text{Attention}(\mathcal{V}^{(l)},\tilde{h}^{(l-1)})\big{)},$ (11)
where $\tilde{h}^{(l-1)}=R({h}^{(l-1)})$. Through this attention mechanism, we expect the learnable vectors $\mathcal{V}^{(l)}$ and the augmented feature representations $\tilde{h}^{(l-1)}$ to cooperatively optimize and enhance robustness against adversarial attacks. Consequently, the final learning objective can be expressed as an extension of Equation (8):
$\min_{\begin{subarray}{c}\mathbf{V}_{i},\,\mathbf{V}_{t},\\ \theta_{i},\,\theta_{t}\end{subarray}}\mathcal{L}(x_{adv},y)$ $=\max_{\begin{subarray}{c}\mathbf{V}_{i},\,\mathbf{V}_{t},\\ \theta_{i},\,\theta_{t}\end{subarray}}s((\mathbf{V}_{i},x_{adv};R_{i}),(\mathbf{V}_{t},y;R_{t})),$ (12)
where $\theta_{i}$ and $\theta_{t}$ represent the learnable parameters of the TokenRefiner $R$ in the visual and text branches, respectively.
The Neural Augmentor implements a feature purification mechanism that operates across three primary dimensions. At the token level, each representation is individually refined, allowing for localized corrections that address token-specific perturbations. At the layer level, TokenRefiner deployed at different network depths learn to address distinct types of perturbations, from low-level feature distortions to high-level semantic shifts. At the modality level, separate TokenRefiner for text and visual pathways allow for specialized correction patterns that address the unique vulnerabilities of each modality.
This multi-dimensional purification approach enables a comprehensive defense against diverse adversarial attacks, addressing perturbations at their source—the feature representations themselves. Rather than treating adversarial robustness as a classification boundary problem, our approach directly targets the mechanism by which adversarial examples cause misclassifications: the distortion of feature representations. By restoring these representations to their clean counterparts, we enable the model to maintain its performance even in the presence of adversarial perturbations.
IV-D Training Methodology
The efficacy of our NAP-Tuning framework relies on a theoretically grounded training methodology that addresses the fundamental challenge in adversarial learning: optimizing the trade-off between standard accuracy and adversarial robustness while maintaining feature-space integrity. Unlike conventional adversarial training approaches that focus primarily on decision boundary robustification, our method directly targets feature-level purification—a strategy particularly well-aligned with the core operating principle of VLMs, which fundamentally depends on cross-modal feature alignment.
IV-D1 Adversarial Example Generation
During training, we generate adversarial examples using the Projected Gradient Descent (PGD) method with the following objective:
$x_{adv}=\arg\max_{x^{\prime}}\mathcal{L}_{adv}(f_{\theta}(x^{\prime}),y)\quad\text{s.t.}\quad\|x^{\prime}-x\|_{\infty}\leq\epsilon,$ (13)
where $f_{\theta}$ represents our model and $\mathcal{L}_{adv}$ denotes the cross-entropy loss. We employ a standard multi-step PGD implementation to generate strong adversarial examples, ensuring that our defense is trained against sophisticated perturbations.
Crucially, the feature-level corrections learned by our model exhibit transferability across different attack types due to the commonality in how various attacks distort the underlying feature manifold. This transferability represents a significant advantage of our feature purification approach over methods that merely harden decision boundaries against specific attack patterns.
IV-D2 Principled Loss Formulation
In contrast to previous defense methods that employ complex, multi-term loss functions with heuristically determined components, we derive a principled, minimal training objective from the theoretical foundations of robust learning:
$\mathcal{L}(\theta)=\mathcal{L}_{clean}(\theta)+\alpha(\tau)\cdot\mathcal{L}_{adv}(\theta),$ (14)
where $\alpha(\tau)$ is a theoretically motivated dynamic balancing coefficient that evolves with training epoch $\tau$. The clean loss $\mathcal{L}_{clean}$ and adversarial loss $\mathcal{L}_{adv}$ are defined as:
$\mathcal{L}_{clean}(\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}(f_{\theta}(x),y)],$ (15)
$\mathcal{L}_{adv}(\theta)$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}(f_{\theta}(x_{adv}),y)],$ (16)
where $\mathcal{D}$ represents the data distribution, and expectations are taken over the training dataset.
The dynamic balancing coefficient $\alpha(\tau)$ follows a sigmoid schedule that systematically transitions from emphasizing clean performance to prioritizing adversarial robustness:
$\alpha(\tau)=\alpha_{0}\cdot\frac{1}{1+\exp\left(-10\left(\frac{\tau}{T_{max}}-0.5\right)\right)},$ (17)
where $\alpha_{0}$ is the maximum weight assigned to the adversarial loss, and $T_{max}$ denotes the total number of training epochs. This schedule implements a curriculum learning strategy that allows the model to first establish representational capacity on clean data before gradually adapting to adversarial inputs.
The theoretical justification for this scheduling approach stems from optimization landscape analysis: the loss surface for adversarial examples typically contains sharper curvature and more local minima than that of clean examples. By initially focusing on clean examples, we guide optimization toward regions of the parameter space with favorable generalization properties before refining these parameters to accommodate adversarial inputs. This procedure effectively navigates the complex optimization landscape of robust learning while mitigating the well-documented trade-off between standard accuracy and adversarial robustness.
TABLE I: Evaluation results on ViT-B16 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method. Best results are highlighted in bold.
ImageNet
Caltech
DTD
Eurosat
Aircraft
Food
Flowers
Pets
Cars
SUN
UCF
Avg
Clean Vanilla $66.7$ $93.3$ $44.1$ $48.3$ $24.7$ $85.8$ $70.7$ $89.1$ $65.6$ $62.6$ $67.5$ $65.3$
FAP $59.7$ $92.2$ $60.8$ $67.8$ $25.4$ $71.3$ $84.5$ $87.1$ $61.1$ $64.0$ $75.3$ $70.6$
AdvPT $67.5$ $94.1$ $69.6$ $69.0$ $28.7$ $85.0$ $87.9$ $91.7$ $70.8$ $70.7$ $77.4$ $73.8$
AdvPT-V $60.7$ $89.8$ $36.9$ $28.1$ $16.1$ $63.8$ $54.4$ $84.7$ $51.7$ $58.4$ $54.1$ $57.0$
AdvPT-VLI $63.8$ $85.7$ $37.2$ $18.2$ $12.0$ $62.3$ $53.6$ $78.7$ $46.9$ $52.8$ $50.5$ $55.4$
AdvPT-VLJ $65.9$ $88.0$ $37.8$ $20.0$ $10.3$ $69.6$ $59.7$ $82.1$ $53.5$ $56.6$ $55.8$ $60.0$
[…]
CWA Vanilla $50.1$ $91.5$ $40.5$ $18.3$ $21.9$ $76.1$ $66.7$ $85.9$ $60.9$ $59.8$ $61.8$ $61.1$
FAP $49.6$ $91.8$ $56.8$ $23.2$ $24.3$ $66.6$ $82.2$ $85.4$ $58.6$ $61.2$ $70.6$ $65.6$
AdvPT $52.1$ $93.1$ $61.8$ $46.9$ $30.8$ $74.9$ $88.2$ $89.8$ $67.8$ $70.5$ $74.3$ $73.0$
AdvPT-V $51.2$ $88.4$ $36.0$ $23.5$ $15.4$ $56.5$ $53.3$ $81.9$ $48.3$ $55.8$ $51.6$ $51.3$
AdvPT-VLI $54.4$ $84.7$ $35.4$ $17.7$ $11.4$ $55.5$ $52.5$ $76.9$ $45.2$ $50.1$ $48.5$ $50.8$
AdvPT-VLJ $53.0$ $86.5$ $35.2$ $16.4$ $8.5$ $52.6$ $50.7$ $77.7$ $42.5$ $51.0$ $48.6$ $49.1$
[…]
AGS Vanilla $53.5$ $86.8$ $30.9$ $11.0$ $18.1$ $56.9$ $56.9$ $78.0$ $54.1$ $53.7$ $49.7$ $54.3$
FAP $51.1$ $88.2$ $50.2$ $12.3$ $22.2$ $58.9$ $77.4$ $80.1$ $53.4$ $55.0$ $61.3$ $58.5$
AdvPT $54.6$ $88.5$ $51.5$ $26.1$ $26.1$ $55.9$ $76.0$ $80.4$ $62.6$ $62.8$ $64.1$ $61.5$
AdvPT-V $53.0$ $85.8$ $33.2$ $15.6$ $13.6$ $47.6$ $47.7$ $78.2$ $44.3$ $51.3$ $44.3$ $50.8$
AdvPT-VLI $56.3$ $81.1$ $33.6$ $14.7$ $10.6$ $46.2$ $48.5$ $74.3$ $41.5$ $46.2$ $41.7$ $47.5$
AdvPT-VLJ $54.9$ $83.0$ $31.6$ $12.1$ $7.5$ $44.0$ $45.3$ $73.9$ $38.6$ $46.2$ $41.4$ $47.3$
[…]
PGD Vanilla $2.7$ $28.3$ $2.0$ $0.1$ $0.0$ $14.4$ $3.2$ $8.8$ $1.4$ $1.4$ $4.0$ $6.7$
FAP $24.0$ $62.4$ $26.7$ $0.0$ $5.2$ $21.8$ $51.7$ $33.7$ $19.5$ $25.4$ $10.9$ $23.3$
AdvPT $1.5$ $27.0$ $6.6$ $0.2$ $0.6$ $0.9$ $4.3$ $2.8$ $0.7$ $1.6$ $2.1$ $5.2$
AdvPT-V $19.6$ $61.5$ $18.6$ $8.1$ $3.8$ $12.4$ $25.0$ $38.0$ $9.8$ $17.3$ $17.0$ $20.4$
AdvPT-VLI $21.8$ $59.1$ $18.6$ $10.0$ $2.9$ $12.9$ $21.7$ $38.2$ $10.0$ $15.7$ $14.2$ $20.2$
AdvPT-VLJ $20.0$ $58.7$ $16.4$ $10.2$ $2.3$ $11.4$ $20.5$ $32.8$ $8.6$ $14.9$ $12.7$ $18.9$
[…]
AutoAttack Vanilla $0.0$ $0.0$ $0.1$ $0.1$ $0.1$ $0.0$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$
FAP $1.6$ $1.1$ $4.7$ $2.3$ $2.7$ $1.4$ $1.1$ $1.9$ $1.4$ $2.6$ $1.8$ $2.2$
AdvPT $0.1$ $0.0$ $0.2$ $0.2$ $0.4$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.0$ $0.1$
AdvPT-V $14.9$ $55.3$ $15.0$ $2.2$ $2.2$ $8.6$ $18.8$ $32.1$ $5.9$ $12.5$ $13.0$ $17.9$
AdvPT-VLI $17.0$ $54.3$ $14.3$ $9.2$ $1.6$ $9.1$ $16.2$ $33.6$ $4.8$ $11.4$ $11.3$ $17.0$
AdvPT-VLJ $10.3$ $43.9$ $10.3$ $3.7$ $1.2$ $4.4$ $9.7$ $19.8$ $1.9$ $6.7$ $6.0$ $11.1$
[…]
TABLE II: Evaluation results on ViT-B32 under clean and adversarial settings. Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method. Best results are highlighted in bold.
ImageNet
Caltech
DTD
Eurosat
Aircraft
Food101
Flowers
Pets
Cars
SUN
UCF
Avg
[…]
CWA Vanilla $48.6$ $90.5$ $40.6$ $33.4$ $16.1$ 74.6 $61.8$ $85.2$ $55.7$ $59.5$ $60.1$ $60.1$
FAP $50.8$ $90.4$ $55.4$ $40.1$ $21.7$ $63.9$ $81.2$ $80.6$ $48.9$ $61.5$ $62.1$ $59.7$
AdvPT 51.2 92.9 60.0 $54.9$ $26.7$ $73.7$ $88.2$ 87.5 $65.5$ 69.3 74.3 $67.7$
AdvPT-V $48.0$ $89.2$ $36.5$ $10.8$ $14.3$ $59.0$ $52.5$ $78.0$ $43.3$ $54.3$ $50.5$ $48.8$
AdvPT-VLI $50.7$ $87.5$ $29.4$ $23.5$ $8.4$ $57.9$ $49.2$ $80.2$ $38.3$ $48.7$ $48.3$ $47.5$
AdvPT-VLJ 51.2 $83.9$ $29.1$ $21.7$ $7.4$ $56.7$ $46.0$ $80.2$ $35.7$ $49.3$ $48.6$ $46.3$
[…]
AGS Vanilla $46.1$ $85.0$ $29.8$ $11.8$ $13.4$ $50.9$ $49.1$ $73.6$ $47.5$ $49.6$ $46.9$ $45.8$
FAP 52.0 89.9 54.4 $31.6$ $21.7$ 63.4 $81.0$ 80.1 $48.8$ 61.1 $61.5$ $58.7$
AdvPT $48.6$ $88.0$ $48.5$ $29.9$ $23.9$ $50.5$ $71.9$ $74.9$ $57.3$ $57.5$ $60.1$ $55.6$
AdvPT-V $47.7$ $85.9$ $33.6$ $3.6$ $12.3$ $49.3$ $46.7$ $73.6$ $37.9$ $48.9$ $44.3$ $44.0$
AdvPT-VLI $50.6$ $83.8$ $26.4$ $14.7$ $6.7$ $49.3$ $43.0$ $76.1$ $33.9$ $43.7$ $41.8$ $42.7$
AdvPT-VLJ $51.2$ $80.0$ $26.7$ $14.6$ $7.2$ $48.0$ $39.5$ $74.5$ $32.0$ $44.5$ $43.1$ $41.9$
[…]
PGD Vanilla $2.0$ $29.1$ $4.9$ $0.2$ $0.0$ $7.4$ $1.7$ $4.2$ $0.4$ $1.5$ $3.0$ $6.7$
FAP $23.4$ $67.1$ $32.8$ $0.4$ $7.0$ $26.2$ $57.3$ $40.0$ $17.6$ $29.8$ $31.5$ $23.3$
AdvPT $2.8$ $31.7$ $10.0$ $0.4$ $0.8$ $1.8$ $8.7$ $3.2$ $2.0$ $2.9$ $4.5$ $5.2$
AdvPT-V $19.0$ $64.6$ $19.8$ $0.3$ $2.9$ $15.0$ $23.0$ $36.3$ $9.6$ $18.9$ $18.3$ $20.4$
AdvPT-VLI $20.7$ $62.1$ $16.3$ $5.7$ $1.5$ $15.9$ $19.3$ $37.6$ $8.4$ $16.2$ $17.7$ $20.2$
AdvPT-VLJ $20.9$ $59.9$ $15.2$ $10.1$ $2.0$ $15.3$ $18.8$ $37.5$ $9.6$ $16.9$ $18.2$ $18.9$
[…]
AutoAttack Vanilla $0.0$ $0.4$ $0.3$ $0.1$ $0.1$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.1$ $0.1$
FAP $2.0$ $1.3$ $5.0$ $1.6$ $3.5$ $2.0$ $1.9$ $1.5$ $1.0$ $3.2$ $2.3$ $2.2$
AdvPT $0.0$ $0.6$ $0.5$ $0.1$ $0.2$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$ $0.2$
AdvPT-V $8.3$ $44.8$ $12.4$ $0.0$ $0.5$ $5.4$ $8.8$ $16.3$ $2.8$ $7.4$ $7.2$ $10.0$
AdvPT-VLI $9.3$ $44.2$ $10.5$ $0.2$ $0.4$ $5.6$ $7.6$ $16.4$ $2.1$ $6.2$ $6.8$ $10.1$
AdvPT-VLJ $10.1$ $44.7$ $10.3$ $1.2$ $0.6$ $6.1$ $7.9$ $18.9$ $2.3$ $7.2$ $8.3$ $11.0$
[…]
V Experiments
This section systematically evaluates the effectiveness of our proposed NAP-Tuning. We first introduce the experimental setup, then validate the key components of our method through a series of carefully designed experiments, and comprehensively compare with existing approaches.
V-A Experimental Setup
V-A1 Datasets and Models
To comprehensively evaluate the effectiveness of our approach, we conduct tests on 11 widely used image classification datasets, including ImageNet [21], Caltech101 [22], DTD [23], EuroSAT [24], FGVC Aircraft [25], Food101 [26], Oxford Flowers [27], Oxford Pets [28], Stanford Cars [29], SUN397 [30], and UCF101 [31]. These datasets encompass a diverse range of visual tasks from fine-grained recognition to scene classification. We follow the training and testing splits defined in [7]. For the ImageNet test set, consistent with prior adversarial attack studies [32, 33, 34], we randomly sample 1,000 images, ensuring one image per class.
Our experiments are conducted on the CLIP model, where the default configuration employs the publicly available ViT-B/16 architecture unless otherwise specified. We also include results for the ViT-B/32 model as a supplementary evaluation. In accordance with previous work, we use hand-crafted prompts as textual inputs (e.g., “a photo of a <class>, a type of pet” for Oxford Pets).
V-A2 Attack Methods
To evaluate adversarial robustness, we implement both white-box and black-box adversarial attacks with more stringent settings compared to our conference version. For white-box attacks, we use learnable textual vectors to construct cross-entropy loss and employ stronger PGD-100 [4] and AutoAttack [35]. For black-box attacks, we utilize more advanced methods such as CWA [36] and AGS [37].
V-A3 Defense Baselines
We compare our method against several strong baselines, including variants of AdvPT and the recent FAP [5]:
  • AdvPT: the original Adversarial Prompt Tuning method, which performs prompt tuning exclusively on the text side. It adapts the textual input prompts to improve robustness without modifying the visual inputs.
  • AdvPT-V: an extension of AdvPT that applies prompt tuning on the vision side. Specifically, it introduces additional learnable prompts into the visual encoder, enabling adaptation of the visual input space to adversarial perturbations.
  • AdvPT-VLJ (vision-language joint prompt tuning): this variant simultaneously introduces learnable prompts to both the vision and text modalities. Moreover, it establishes a mapping between the vision and language prompts, allowing joint adversarial tuning of both modalities. This structure follows a design similar to MaPLe [8].
  • AdvPT-VLI (vision-language independent prompt tuning): a modification of AdvPT-VLJ that removes the mapping between vision and language prompts. In this setup, the learnable prompts for the visual and textual inputs are optimized independently, without explicit cross-modal interactions.
  • FAP [5]: a recent method proposed for few-shot adversarial robustness, which carefully designs the objective function.
V-A4 Implementation Details
Our training framework spans 90 epochs with a batch size of 512. We adopt the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is modulated through a cosine annealing schedule. For adversarial example generation during training (as defined in Equation (12)), we implement PGD-5 with a default perturbation magnitude $\epsilon=1/255$ for both training and testing phases, though we analyze the effects of different $\epsilon$ values in subsequent sections. For the TokenRefiner $R$, we implement a default 2-layer neural network architecture, with the impact of network depth examined in our ablation studies. Regarding the multi-layer prompt architecture, our default configuration utilizes all 12 transformer layers, with layer-specific effects analyzed in detailed experiments. The hyperparameter $\alpha_{0}$ in Equation (17) is set to 5.0. All experiments are conducted on NVIDIA A800 80GB GPUs.
V-B Main Results
The main evaluation results are summarized in Tables I and II. Overall, NAP-Tuning achieves highly competitive clean accuracy compared to existing baselines. While its clean performance is slightly lower than that of AdvPT, this is expected, as AdvPT explicitly prioritizes clean accuracy in the robustness-accuracy trade-off, often at the cost of adversarial robustness.
In terms of adversarial robustness, NAP-Tuning demonstrates clear and consistent advantages. Under black-box attacks, particularly CWA and AGS, our method outperforms all baselines across both ViT-B16 and ViT-B32 backbones, highlighting its superior generalization to unseen adversarial threats. More importantly, under white-box attacks such as PGD and AutoAttack, NAP-Tuning exhibits significant improvements. AutoAttack is widely regarded as the most rigorous and comprehensive white-box evaluation benchmark, combining several strong attacks to provide a thorough assessment of model robustness. Under this challenging setup, NAP-Tuning achieves substantial gains over all competitors: on ViT-B16, it attains 51.4% on average under AutoAttack, outperforming the strongest baseline (AdvPT-V) by 33.5%; on ViT-B32, it reaches 44.0%, exceeding the best competitor (AdvPT-VLJ) by 33.0%. These results clearly demonstrate the robustness of NAP-Tuning against powerful, diverse adversarial threats.
V-C Multi-layer Prompt Architecture
To investigate how the depth of our multi-layer prompt architecture affects model robustness, we conduct an ablation study by varying the number of layers where prompts are inserted, from 1 to 12. Fig. 4 presents the performance on eleven diverse datasets under both clean and adversarial conditions.
The results reveal several key insights. First, as shown in the average performance (Fig. 4l), both clean and robust accuracy generally improve as the number of layers increases, confirming the effectiveness of our hierarchical prompting strategy. Interestingly, we observe dataset-specific patterns related to task complexity. For more challenging datasets such as ImageNet (Fig. 4f), SUN397 (Fig. 4j), and Food101 (Fig. 4e), performance follows an inverted U-shape, peaking at intermediate layer depths before declining. This pattern indicates mild overfitting when increasing parameter capacity without corresponding increases in training examples (shot=16). The highlighted regions in these plots mark the optimal layer ranges for these datasets.
These findings reveal the potential for further performance enhancements through larger training sets and dataset-specific layer optimization, suggesting that improved robustness can be achieved through straightforward data augmentation. Nevertheless, our method demonstrates strong performance even with the default configuration (layer=12) as shown in Table I and Table II, consistently outperforming baseline approaches across diverse datasets.
Figure 4: Clean and robust accuracy across datasets when varying the number of prompt layers (1-12). Complex datasets (ImageNet, Food101, SUN397) show optimal performance at intermediate depth, while other datasets benefit from deeper architectures.
V-D Adversarial Regularization Parameter
We analyze the impact of adversarial regularization weight ($\alpha$) on both standard (clean) and adversarial (robust) accuracy. Fig. LABEL:fig:alpha presents results across $\alpha$ values from 0.01 to 100, with panel (a) showing both metrics on a single y-axis and panel (b) using dual y-axes to highlight relative trends.
The results show that clean accuracy remains stable (98.5%-99.1%) across all $\alpha$ values, while robust accuracy increases substantially from 79.9% at $\alpha$=0.01 to 93.9% at $\alpha$=2.0, with diminishing gains thereafter. For $\alpha\geq 5.0$, robust accuracy consistently exceeds 94%. Rather than specifically optimizing for the peak performance at $\alpha$=20.0 (94.4% robust accuracy), we adopt $\alpha$=5.0 as our representative configuration, which achieves 98.8% clean accuracy and 94.2% robust accuracy. These findings demonstrate that our approach effectively balances clean accuracy and adversarial robustness across a broad range of $\alpha$ values, exhibiting strong performance without requiring fine-tuned parameter selection.
V-E TokenRefiner Architecture
We analyze the impact of TokenRefiner network depth on model performance using Oxford Flowers (simpler task) and ImageNet (more challenging task). Fig. LABEL:fig:token presents clean and robust accuracy across different TokenRefiner layer configurations. Results demonstrate that TokenRefiner is essential for model convergence—its absence (layer = 0) prevents the model from converging under 5-step PGD training, resulting in poor performance (4.6% clean accuracy on ImageNet, 25.8% on Oxford Flowers). Performance improves substantially as layers increase from 0 to 2, with ImageNet clean accuracy rising from 4.6% to 54.4% and robust accuracy from 1.9% to 28.0%. Similarly, Oxford Flowers shows improvements from 25.8% to 96.1% (clean) and from 15.1% to 86.6% (robust). Performance stabilizes beyond 2 layers, leading us to select a 2-layer architecture for our final implementation. Notably, unlike previous approaches such as FAP [5] and APD [6], which lack specialized prompt structure designs and are therefore limited to weaker adversarial search strategies, our TokenRefiner enables convergence under stronger adversarial perturbations. The consistent pattern across datasets of varying complexity suggests our approach generalizes well without requiring task-specific architectural adjustments.
V-F Perturbation Magnitude
We analyze the impact of perturbation budgets during both training and testing phases. Table III shows results on the Flowers dataset with varying $\epsilon$ values. The results reveal a clear trade-off between clean accuracy and adversarial robustness. As training perturbation budget increases from $\epsilon=1$ to $\epsilon=4$, clean accuracy decreases from 96.1% to 87.5%, while robustness against stronger attacks improves substantially. Models trained with larger $\epsilon$ values perform significantly better when tested under matching or higher perturbation budgets—a model trained with $\epsilon=4$ achieves 55.9% accuracy when tested at the same level, versus only 17.1% for a model trained with $\epsilon=1$. Models experience rapid performance degradation when tested with perturbations larger than their training budget. This observation suggests that, similar to traditional adversarial training, adversarial prompt tuning exhibits a clear trade-off between clean accuracy and adversarial robustness.
TABLE III: Accuracy (%) under $\ell_{\infty}$-bounded adversarial perturbations. Models are trained and evaluated with varying perturbation magnitudes (measured in $\epsilon/255$). Bold values indicate the highest robustness at each test-time perturbation level.
Training $\bm{\epsilon}$ Clean Accuracy Robust Accuracy @ Test-time $\bm{\epsilon}$
1 2 4 8
1 96.1 86.5 67.0 17.1 0.1
2 94.2 87.2 75.3 38.1 1.3
4 87.5 79.2 75.7 55.9 11.8
V-G Few-shot Learning
We examine the impact of shot count on model performance. Fig. 6 illustrates the relationship between number of shots (1-16) and model accuracy across diverse datasets. Clean accuracy shows consistent improvement, with average accuracy increasing from 40.6% (single shot) to 73.2% (16 shots)—a 1.8× improvement. More remarkably, robust accuracy exhibits steeper relative gains, rising from 14.1% to 49.8%—a dramatic 3.5× improvement.
This differential impact reveals a key insight: while few-shot learning is known to enhance clean accuracy, its benefits for robust performance are substantially more pronounced. The results indicate that robustness requires a richer representation of class concepts that becomes increasingly available with additional examples. These findings align with our observations in Section V-C, confirming that increasing shot count offers tremendous potential for performance enhancement in our method, providing a straightforward path to significantly improved robustness without architectural modifications.
Figure 6: Impact of shot count on clean and robust accuracy across eleven datasets. Each subplot (a-k) shows performance on individual datasets, while subplot (l) presents the average across all datasets.
V-H Context Vector Count
We examine how the number of learnable context vectors ($\mathcal{V}_{t}$) affects model performance. Fig. 7 shows results for models with 1-16 context vectors on ImageNet and Oxford Flowers datasets. For ImageNet, clean accuracy remains relatively stable (55.9%-58.1%) across different vector counts, while robust accuracy shows modest improvements with increased vectors, peaking at 29.6% with 14-16 vectors. On Oxford Flowers, both metrics maintain high performance across all configurations, with clean accuracy ranging from 95.5% to 96.5% and robust accuracy between 85.2% and 86.7%. These findings suggest that the number of context vectors $\mathcal{V}_{t}$ has a relatively minor impact on overall performance compared to the influence of layer count in Multi-Layer Prompts observed in Section V-C. This indicates that input-level representational capacity is not the primary performance bottleneck, particularly for less complex datasets where even few context vectors prove sufficient.
Figure 7: Effect of context vector count (1-16) on ImageNet and Oxford Flowers. Performance varies minimally across different vector counts, indicating context vectors are not a primary bottleneck for model capability.
VI Conclusion and Discussion
This paper presents the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning), a novel architecture-focused approach to enhancing adversarial robustness in Vision-Language Models. Our work addresses significant limitations in existing adversarial prompt tuning methods by fundamentally reconceptualizing prompt architecture through feature purification mechanisms.
VI-A Conclusion
We have extended adversarial prompt tuning from text-only to multi-modal contexts and from single-layer to multi-layer architectures while introducing token refiners that enable modality-specific and layer-specific feature correction. This architectural innovation represents a paradigm shift from previous loss-focused approaches, emphasizing the critical role of feature purification in addressing adversarial distortions. Through extensive experimentation, we have demonstrated that NAP-Tuning substantially outperforms existing methods, achieving state-of-the-art robustness against strong attacks while maintaining competitive clean accuracy.
VI-B Broader Implications
Our findings have several significant implications: 1) highlight the insufficiency of merely transplanting prompt tuning techniques from generalization domains to adversarial defense without architectural reconsideration; 2) underscore the importance of addressing adversarial perturbations at the feature level rather than solely through loss modification.
VI-C Discussion
Future work could explore adaptive token refiners that dynamically adjust their behavior based on detected perturbation characteristics, integration of uncertainty quantification mechanisms to improve refinement reliability, and extension to other multi-modal tasks beyond classification. Additionally, investigating the theoretical connections between feature purification and information bottleneck principles could yield deeper insights into adversarial robustness in representation learning.
In conclusion, our work establishes that addressing adversarial robustness in VLMs requires fundamental architectural innovations rather than merely adapting existing prompt tuning techniques. The Neural Augmentor framework provides a principled approach to this challenge, offering both improved performance and conceptual insights that advance our understanding of robust multi-modal learning.

### v2

@@
 Building upon our previous work on Adversarial Prompt Tuning (AdvPT), which introduced learnable text prompts to enhance adversarial robustness in VLMs without extensive parameter training, we present a significant extension by introducing the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning).
-Our key innovations include: (1) extending AdvPT from text-only to multi-modal prompting across both text and visual modalities, (2) expanding from single-layer to multi-layer prompt architectures, and (3) proposing a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification to directly address the distortions introduced by adversarial attacks in feature space.
-Our NAP-Tuning approach incorporates token refiners that learn to reconstruct purified features through residual connections, allowing for modality-specific and layer-specific feature correction.
+As a significant extension, NAP-Tuning first establishes a comprehensive multi-modal (text and visual) and multi-layer prompting framework.
+The core of this framework is a targeted structural augmentation for feature-level purification, implemented through our Neural Augmentor approach.
+This framework implements feature purification by incorporating TokenRefiners—lightweight neural modules that learn to reconstruct purified features via residual connections—to directly address distortions in the feature space.
+This structural intervention is what enables the multi-modal and multi-layer system to effectively perform modality-specific and layer-specific feature rectification.
 Comprehensive experiments demonstrate that NAP-Tuning significantly outperforms existing methods across various datasets and attack types.
-Notably, our approach shows significant improvements over the strongest baselines under the challenging AutoAttack benchmark, outperforming them by 33.5% on ViT-B16 and 33.0% on ViT-B32 architectures while maintaining competitive clean accuracy.
-This work establishes the importance of architecture redesign in prompt tuning for adversarial robustness, moving beyond loss-focused approaches to create an adaptive defense mechanism that can identify and rectify adversarial perturbations across embedding spaces.
-Index Terms:
-Adversarial robustness, Vision-Language models, Prompt tuning, Feature purification, Neural augmentation
+Notably, our approach shows significant improvements over the strongest baselines under the challenging AutoAttack benchmark, outperforming them by 32.3% on ViT-B16 and 31.3% on ViT-B32 architectures while maintaining competitive clean accuracy.
+This work highlights the efficacy of internal feature-level intervention in prompt tuning for adversarial robustness, moving beyond input-side alignment approaches to create an adaptive defense mechanism that can identify and rectify adversarial perturbations across embedding spaces.
 I Introduction
@@
 Following our initial work, several approaches [41, 15] have attempted to address some of these issues, particularly by extending prompting to multiple modalities and layers.
-However, these incremental improvements fail to address the core architectural challenge of prompt tuning in adversarial settings.
-The fundamental limitation across all existing approaches—including both our original AdvPT and subsequent works—is the direct transplantation of prompt tuning techniques from the generalization domain (e.g., CoOp [7] and MaPLe [8]) to adversarial defense without rethinking the underlying architecture.
-This approach overlooks a critical insight: prompt tuning, originally developed for clean data environments, faces fundamentally different challenges in adversarial scenarios where feature spaces are deliberately corrupted.
-Adversarial perturbations introduce systematic distortions that cannot be effectively countered through loss modifications alone, regardless of whether the prompting occurs in single or multiple modalities, or at one or multiple layers.
-To comprehensively address all three limitations, we propose the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning).
-Our approach not only extends prompting to multiple modalities and layers but, most critically, fundamentally reconceptualizes prompt architecture for adversarial defense by integrating feature purification mechanisms.
-The Neural Augmentor framework incorporates token refiners—lightweight neural networks that learn to reconstruct clean feature representations from adversarially perturbed inputs through residual connections.
-This architecture enables modality-specific and layer-specific feature correction, yielding substantially enhanced adversarial robustness.
+However, these works, while incrementally beneficial, share the same fundamental architectural limitations as AdvPT and thus fail to address the core challenge of robust feature representation.
+The fundamental limitation across all existing prompt-based approaches is the direct transplantation of prompt tuning techniques from the generalization domain (e.g., CoOp [40] and MaPLe [8]) to adversarial defense without rethinking the underlying architecture.
+This overlooks a critical insight: adversarial robustness is a fundamentally harder task than standard generalization, as it requires richer model capacity [25, 36].
+This requirement for increased capacity creates a fundamental trade-off during VLM adaptation.
+While freezing the backbone preserves pre-trained knowledge, the limited structural capacity of standard prompt tuning, which typically optimizes a small set of input-side vectors, often fails to rectify corrupted features effectively.
+Conversely, high-capacity full fine-tuning risks catastrophic forgetting of pre-trained representations [28].
+We aim to bridge this gap by incorporating a modular structural augmentation that provides targeted capacity for feature purification while maintaining the backbone’s integrity.
+A systematic analysis and positioning of these adaptation paradigms are further detailed in Section II-D.
+To address these limitations, we propose the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning).
+Our approach introduces a targeted structural augmentation designed to navigate the trade-off between model capacity and knowledge preservation.
+We keep the backbone frozen to avoid catastrophic forgetting, but introduce a modular neural component: the Neural Augmentor.
+Unlike standard prompt vectors, the Augmentor is a set of lightweight neural networks (TokenRefiners) that operate directly on the feature representations.
+This modular design provides the functionality required to model and rectify feature distortions at the token level.
 Fig.
-LABEL:fig:comparison illustrates the evolution from text-only prompting to our comprehensive NAP-Tuning, demonstrating how our method facilitates clean feature reconstruction from adversarially corrupted inputs across both modalities and multiple network layers.
+3(a) illustrates the evolution from text-only prompting to our comprehensive NAP-Tuning.
+(a) Comparison of adversarial prompt tuning approaches: (a) The original AdvPT that uses only text prompts, (b) Existing multimodal approaches that incorporate prompts in both visual and text pathways, and (c) Our proposed NAP-Tuning framework that extends multimodal prompting with feature purification via token refiners, enabling the reconstruction of clean feature representations from adversarially perturbed inputs.
 Through extensive experiments under rigorous evaluation protocols, we demonstrate that our approach significantly outperforms our initial AdvPT and other state-of-the-art methods, achieving superior robustness while maintaining competitive clean accuracy.
 Our analysis revealed that NAP-Tuning effectively learns to correct adversarial distortions in feature space, providing insights into the working mechanisms of successful adversarial defenses.
-As an extension of our conference paper [3], this work makes several significant contributions:
-  • Multi-modal defense framework: We extend the initial text-only AdvPT to a comprehensive multi-modal framework that incorporates learnable prompts in both text and visual pathways, creating a more effective defense mechanism that addresses vulnerabilities across modalities.
-  • Multi-layer prompt architecture: We expand from single-layer prompting to multi-layer prompt architectures, allowing for more nuanced and layer-specific defenses that can address adversarial perturbations manifesting differently at various depths of the network.
-  • Neural Augmentor design: Most significantly, we propose a novel architecture-level redesign through our Neural Augmentor approach, which implements feature purification via token refiners.
-This represents a paradigm shift from previous loss-focused prompt tuning methods to architecture-focused prompt tuning defense, addressing the fundamental limitations of existing approaches.
-  • Comprehensive evaluation: We provide extensive experimental validation across multiple datasets and attack types under stricter evaluation settings.
-Our approach demonstrates significant performance improvements over the previous conference version (AdvPT), achieving average gains of 42.6% under PGD-100 attacks and 47.5% under the more challenging AutoAttack benchmark, establishing new standards for adversarial robustness in VLMs.
+As an extension of our conference paper [37], which introduced the concept of Adversarial Prompt Tuning, this work makes several significant contributions:
+  • Systematic Extension: Building upon our original text-only AdvPT [37], we implement a coordinated framework across both visual and textual pathways.
+This systemic extension addresses adversarial vulnerabilities that manifest across modalities, providing a unified multi-layer prompting architecture for robust vision-language adaptation.
+  • Modular Structural Augmentation: We introduce the Neural Augmentor, a modular structural augmentation utilizing lightweight TokenRefiners for explicit feature-level purification.
+This design represents a functional shift from loss-alignment methods to direct feature rectification, providing a plug-and-play mechanism that is easily compatible with various frozen VLM backbones.
+  • Comprehensive Validation: We provide extensive experimental validation for our method across 11 datasets and under stricter evaluation settings.
+Our analysis confirms the effectiveness of the feature purification mechanism, demonstrating state-of-the-art robustness gains while preserving generalization.
+TABLE I: Comparative analysis of Adversarial Adaptation Paradigms for VLMs. We analyze different strategies based on their architectural approach to the robustness-generalization trade-off.
+Adaptation Paradigm (Examples) Preserves Prior Increases Model Architectural Trainable Resulting Robustness
+Knowledge? Capacity? Redesign? Parameters Potential
+(No Adaptation)
+Full Fine-Tuning (TeCoA [18]) N/A a High Low
+Prompt Tuning (FAP [41], APD [15]) Low Low Medium
+Neural Augmentation (Ours) High Low-Medium High
+  • a
+    This paradigm does not add new capacity, but unlocks all existing (Very High) parameters for training.
 II Related Work
 II-A Vision-Language Models
-VLMs have established themselves as powerful frameworks for bridging visual and textual modalities.
-These models can be categorized into two principal architectures.
-The first category comprises generative models built upon Large Language Models (LLMs), such as LLaVA [9] and Qwen-VL [10], which extend text-only LLMs with vision encoders to enable multimodal understanding and generation.
-The second category—which is the focus of adversarial prompt tuning techniques—utilizes contrastive learning to establish joint embeddings of images and text.
-Models such as CLIP [1] and ALIGN [2] project both modalities into a unified semantic space where related content exhibits higher similarity.
-These models demonstrate exceptional capabilities across diverse tasks through their dual-encoder architecture.
-However, their susceptibility to adversarial attacks, particularly in the visual domain, poses significant security concerns as their deployment expands into critical applications.
+VLMs establish a semantic bridge between visual and textual modalities, typically following two architectural paradigms: generative models based on Large Language Models (e.g., LLaVA [13], Qwen-VL [35]) and contrastive models (e.g., CLIP [23], ALIGN [7]) that learn joint embedding spaces.
+While the latter demonstrates exceptional zero-shot transfer capabilities through unified multimodal projections, their inherent susceptibility to adversarial perturbations—particularly in the visual domain—poses significant security risks.
+Ensuring the robustness of these contrastive embeddings is critical as their deployment expands into safety-critical applications.
 II-B Prompt Learning
-Prompt learning emerged in natural language processing as an efficient parameter-efficient adaptation technique for pre-trained language models [11, 12].
-This paradigm focuses on optimizing input transformations rather than modifying model parameters.
-For VLMs, text-based prompt learning methods such as CoOp [7] and CoCoOp [13] demonstrated that replacing hand-crafted templates with learnable context vectors could significantly improve model performance.
-Our previous work, AdvPT [3], applied this concept to adversarial defense, showing that learnable text prompts could enhance robustness against adversarial attacks without modifying the underlying model.
-Multi-modal prompt learning approaches including MaPLe [8] and PromptKD [14] subsequently extended this paradigm by incorporating prompts in both textual and visual pathways.
-These architectural innovations have provided foundational techniques for multi-modal adversarial prompt tuning, though their primary focus on generalization rather than robustness presents limitations when applied to adversarial defense.
+Originating as a parameter-efficient alternative to full fine-tuning in NLP [14, 10], prompt learning adapts pre-trained models by optimizing input-side transformations.
+In the VLM context, methods such as CoOp [40] and CoCoOp [39] replaced manual templates with learnable context vectors.
+Subsequent innovations, including MaPLe [8] and PromptKD [12], extended this paradigm to multi-modal pathways.
+However, as these techniques were primarily engineered for task generalization, their fixed-architecture design presents a fundamental bottleneck when attempting to counteract the structural feature distortions introduced by adversarial attacks.
 II-C Adversarial Prompt Tuning
-Conventional adversarial defense methods such as adversarial training [4], TRADES [15], and MART [16] have proven effective but remain computationally prohibitive for large-scale VLMs.
-These approaches typically require training the whole model parameters from scratch, making them impractical for deployment in VLMs.
-An innovative alternative approach was introduced by Mao et al.
-[17] through their TeCoA framework, which intelligently leverages pre-trained VLM weights as initialization and selectively finetunes CLIP’s vision encoder.
-Our previous work, AdvPT [3], introduced a more efficient and more effective paradigm by applying prompt learning to adversarial defense in VLMs.
-By aligning text embeddings with adversarial image embeddings through learnable prompts, AdvPT demonstrated that robust representations could be achieved with minimal computational overhead.
-Recent extensions to this approach include FAP [5], which leverages limited data to learn adversarially correlated text supervision while enhancing multi-modal feature consistency and uni-modal feature differentiation between natural and adversarial examples.
-Recent advances in test-time optimization have yielded particularly impressive results in defending VLMs against adversarial attacks.
-Sheng et al.
-[18] developed R-TPT, an elegant test-time prompt tuning method that enhances the robustness of vision-language models by minimizing pointwise entropy and introducing a reliability-based weighted ensembling strategy.
-Their approach demonstrates the power of adaptive defenses that operate at inference time.
-TAPT [19] introduces an unsupervised approach that optimizes bimodal prompts at test time to enhance CLIP’s robustness while maintaining clean performance.
-In parallel, Xing et al.
-[20] introduced TTC, an innovative training-free defense mechanism that ingeniously leverages CLIP’s pre-trained vision encoder to counteract adversarial images during inference.
-Their approach offers complementary protection that can be effectively combined with existing robust models, demonstrating the value of ensemble strategies in comprehensive defense frameworks.
-While these methods have achieved incremental improvements through loss function innovations, they overlook a critical consideration: the fundamental mismatch between prompt architectures designed for generalization and the unique requirements of adversarial robustness.
-Addressing adversarial vulnerabilities requires mechanisms specifically engineered to counteract feature-level distortions—a challenge that far exceeds general-purpose representation enhancement.
-This insight motivates our current work, which demonstrates that architectural innovations in prompt design can substantially improve adversarial robustness beyond what is possible through loss function modifications alone.
-By reimagining prompt architecture specifically for adversarial defense, we address the feature distortion mechanisms that underlie successful attacks on VLMs.
+To circumvent the prohibitive computational costs of conventional adversarial training [16, 36, 31], AdvPT [37] introduced an efficient defense paradigm by aligning text prompts with adversarial image embeddings.
+This prompted several extensions, such as FAP [41], which focuses on few-shot adversarial consistency, and APD [15], which leverages bimodal distillation for robust adaptation.
+In parallel, test-time optimization strategies (e.g., R-TPT [26], TAPT [30], TTC [34]) have enhanced VLM security during inference.
+Despite these advancements, existing methods predominantly focus on loss-level alignment of input-side prompt vectors.
+Our work diverges from these trends by introducing a modular structural augmentation designed specifically for internal feature-level rectification, a mechanism that targets the root cause of feature distortion beyond simple distribution alignment.
+II-D Architectural Trade-offs in Robust VLM Adaptation
+The theoretical foundation for adapting VLMs for robustness involves navigating the fundamental trade-off between standard accuracy and adversarial robustness [36].
+Robust learning theory establishes that achieving adversarial generalization is a more difficult task, requiring significantly higher model capacity than standard generalization [25].
+In the context of large pre-trained models, this requirement creates a specific dilemma: while full fine-tuning provides the necessary capacity to learn robust features, it risks the degradation of pre-trained knowledge [28, 38].
+Conversely, fixed-architecture prompt tuning preserves generalization but lacks the structural complexity needed to counteract feature-level distortions.
+We clarify this landscape in Table I, positioning our work within three dominant adaptation paradigms.
+Unlike conventional prompt tuning that relies on loss-alignment of input-side vectors, our Neural Augmentation paradigm introduces a modular structural intervention.
+By incorporating internal TokenRefiners, we provide the targeted capacity required for feature purification without compromising the backbone’s integrity.
+This design is fundamentally modular, allowing the rectification mechanism to be integrated into various frozen transformer backbones as a generic plug-in.
 III Preliminary: Adversarial Prompt Tuning
-To provide context for our extended approach, we first revisit the key concepts of Adversarial Prompt Tuning (AdvPT) as introduced in our previous work [3].
+To provide context for our extended approach, we first revisit the key concepts of Adversarial Prompt Tuning as introduced in our previous work [37].
 III-A Revisiting CLIP
@@
 For an image $x$ and a text description $t$, CLIP computes embeddings $f_{I}(x)$ and $f_{T}(t)$, and the similarity is calculated as:
-$s(x,t)=\frac{f_{I}(x)\cdot f_{T}(t)}{||f_{I}(x)||\cdot||f_{T}(t)||}$ (1)
+[…]
 For classification tasks, CLIP computes the similarity between an image and a set of text templates describing each class (e.g., “a photo of a [CLASS]”), and selects the class with the highest similarity.
@@
 This stronger adversary can generate more effective attacks by maximizing the cross-entropy loss between image-text pairs (even when the text is represented as learnable vectors):
-$x_{adv}$ $=\arg\max_{x^{\prime}}\mathcal{L}_{CE}(f_{I}(x^{\prime}),f_{T}(t_{y}))$
+$x_{adv}$ $=\arg\max_{x^{\prime}}\mathcal{L}_{CE}(cos(f_{I}(x^{\prime}),f_{T}(\mathbf{T})),y)$
 […]
-where $t_{y}$ is the text template for class $y$, and $\mathcal{L}_{CE}$ is the cross-entropy loss measuring image-text similarity.
+where $\mathbf{T}=\{t_{1},...,t_{k}\}$ is the set of all $k$ class text templates, $y$ is the ground-truth class label, $cos(\cdot,\cdot)$ is the cosine similarity function which produces the logits, and $\mathcal{L}_{CE}$ is the standard cross-entropy loss.
 III-C Basic Adversarial Prompt Tuning
@@
 In AdvPT, the standard text template “a photo of a [CLASS]” is replaced with a template containing learnable context vectors:
-[…]
+$(\mathcal{V}_{t},y)=[\mathcal{V}_{t},\text{[CLASS]}],$ (4)
 where $\mathcal{V}_{t}$ are learnable vectors that are optimized to align with adversarial image embeddings, and $y$ represents class label.
@@
 IV Neural Augmentor for Multi-Modal Adversarial Prompt Tuning
-Adversarial attacks fundamentally operate by introducing carefully crafted perturbations to input data that cause neural networks to make incorrect predictions.
-In the context of VLMs, these perturbations manipulate feature representations in ways that disrupt the crucial alignment between visual and textual information.
-Building upon our preliminary work on AdvPT, we introduce a comprehensive redesign of prompt-based adversarial defenses through our NAP-Tuning framework.
-This section details our approach, its architectural innovations, and theoretical foundations.
+Building upon our preliminary work on AdvPT, we introduce the NAP-Tuning framework as a structural augmentation for prompt-based adversarial defenses.
+This section details our approach, its structural augmentations, and theoretical foundations.
 IV-A Framework Overview
@@
 Fig.
-LABEL:fig:comparison (c) introduces the overall framework, while Fig.
-3 illustrates the architecture of our Neural Augmentor module.
-At its core, our approach aims to defend VLMs against adversarial attacks by addressing their fundamental vulnerability: the distortion of feature representations.
-Unlike previous prompt defenses that focus on loss function modifications or robust classification boundaries, our method directly targets the feature corruption mechanism through specialized architectural components.
-The framework consists of three primary components that work in concert.
-First, a multi-modal prompting system introduces learnable prompt vectors that operate simultaneously in both textual and visual pathways of the model.
-Second, a multi-layer prompt architecture implements hierarchical prompting structures at different depths within the transformer architecture to address adversarial effects at various levels of feature abstraction.
-Third, Neural Augmentor modules serve as specialized neural networks that perform feature purification through token refiners.
+3(a) (c) and Fig.
+3 illustrate the NAP-Tuning framework and its core Neural Augmentor module.
+Conceptually, our approach represents a strategic evolution of the efficiency-oriented paradigm established by AdvPT [37].
+While the original AdvPT focused on minimizing computational overhead by avoiding backbone backpropagation, NAP-Tuning identifies a superior trade-off: by incorporating lightweight structural augmentations for internal feature purification, we achieve a massive leap in robustness while maintaining an efficiency profile that remains orders of magnitude superior to traditional full fine-tuning.
+This conceptual shift—moving from purely input-side alignment to internal feature rectification—is implemented through three coordinated components: (1) a multi-modal prompting system for visual-textual coordination, (2) a multi-layer prompt architecture for hierarchical intervention, and (3) Neural Augmentor modules that perform explicit token-level purification via residual refiners.
 IV-B Structural Innovations
@@
 $\hat{f}_{I}(x)\approx f_{I}(x_{adv})+\Phi(f_{I}(x_{adv})),$ (10)
-where $\Phi$ is a learned correction function that approximates $-\Delta f$, effectively “purifying” the adversarial features (including both intermediate and final feature representations) back toward their clean counterparts.
+where $\Phi$ is a learned correction function that approximates $\Delta f$, effectively “purifying” the adversarial features (including both intermediate and final feature representations) back toward their clean counterparts.
 IV-C2 TokenRefiner Architecture
-The core component of the Neural Augmentor is the TokenRefiner $R$—a lightweight neural network that processes individual token representations to identify and correct adversarial distortions.
+The core component of the Neural Augmentor is the TokenRefiner RR—a lightweight neural network that processes individual token representations to identify and correct adversarial distortions.
 For each token representation $z\in\mathbb{R}^{d}$, the TokenRefiner computes a corrective term $\tilde{z}=R(z)$.
@@
 where $\theta_{i}$ and $\theta_{t}$ represent the learnable parameters of the TokenRefiner $R$ in the visual and text branches, respectively.
-The Neural Augmentor implements a feature purification mechanism that operates across three primary dimensions.
-At the token level, each representation is individually refined, allowing for localized corrections that address token-specific perturbations.
-At the layer level, TokenRefiner deployed at different network depths learn to address distinct types of perturbations, from low-level feature distortions to high-level semantic shifts.
-At the modality level, separate TokenRefiner for text and visual pathways allow for specialized correction patterns that address the unique vulnerabilities of each modality.
-This multi-dimensional purification approach enables a comprehensive defense against diverse adversarial attacks, addressing perturbations at their source—the feature representations themselves.
-Rather than treating adversarial robustness as a classification boundary problem, our approach directly targets the mechanism by which adversarial examples cause misclassifications: the distortion of feature representations.
-By restoring these representations to their clean counterparts, we enable the model to maintain its performance even in the presence of adversarial perturbations.
+The Neural Augmentor implements a hierarchical purification mechanism across token, layer, and modality dimensions, enabling localized and structural rectification of adversarial artifacts.
+This approach represents a conceptual shift from treating robustness as a classification boundary problem (typical of input-side alignment methods) to directly targeting the root cause: the distortion of internal feature representations.
+By explicitly intercepting and restoring perturbed tokens within the backbone’s feature flow, our design maintains the model’s representational integrity and cross-modal alignment even under sophisticated adversarial threats.
 IV-D Training Methodology
@@
 This procedure effectively navigates the complex optimization landscape of robust learning while mitigating the well-documented trade-off between standard accuracy and adversarial robustness.
-TABLE I: Evaluation results on ViT-B16 under clean and adversarial settings.
-Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method.
+TABLE II: Evaluation results on ViT-B16 under clean and adversarial settings.
+Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (M-Attack), where $\dagger$ presents our method.
 Best results are highlighted in bold.
@@
 Clean Vanilla $66.7$ $93.3$ $44.1$ $48.3$ $24.7$ 85.8 $70.7$ $89.1$ $65.6$ $62.6$ $67.5$ $65.3$
-FAP $59.7$ $92.2$ $60.8$ $67.8$ $25.4$ $71.3$ $84.5$ $87.1$ $61.1$ $64.0$ $75.3$ $70.6$
-AdvPT $67.5$ $94.1$ $69.6$ $69.0$ $28.7$ $85.0$ $87.9$ $91.7$ $70.8$ $70.7$ $77.4$ $73.8$
-AdvPT-V $60.7$ $89.8$ $36.9$ $28.1$ $16.1$ $63.8$ $54.4$ $84.7$ $51.7$ $58.4$ $54.1$ $57.0$
-AdvPT-VLI $63.8$ $85.7$ $37.2$ $18.2$ $12.0$ $62.3$ $53.6$ $78.7$ $46.9$ $52.8$ $50.5$ $55.4$
-AdvPT-VLJ $65.9$ $88.0$ $37.8$ $20.0$ $10.3$ $69.6$ $59.7$ $82.1$ $53.5$ $56.6$ $55.8$ $60.0$
+FAP $60.1$ $92.8$ $60.7$ $68.3$ $26.0$ $72.5$ $84.6$ $87.7$ $60.9$ $65.2$ $75.1$ $68.5$
+APD $57.7$ $90.2$ $57.3$ $65.7$ $28.9$ $78.2$ $85.7$ $84.7$ $58.8$ $59.6$ $75.8$ $67.5$
+AdvPT 67.5 94.1 69.6 $69.0$ $28.7$ $85.0$ $87.9$ 91.7 $70.8$ 70.7 77.4 73.9
+AdvPT-V $60.7$ $89.8$ $36.9$ $28.1$ $16.1$ $63.8$ $54.4$ $84.7$ $51.7$ $58.4$ $54.1$ $54.4$
+AdvPT-VLI $63.8$ $85.7$ $37.2$ $18.2$ $12.0$ $62.3$ $53.6$ $78.7$ $46.9$ $52.8$ $50.5$ $51.1$
+AdvPT-VLJ $65.9$ $88.0$ $37.8$ $20.0$ $10.3$ $69.6$ $59.7$ $82.1$ $53.5$ $56.6$ $55.8$ $54.5$
 […]
-CWA Vanilla $50.1$ $91.5$ $40.5$ $18.3$ $21.9$ $76.1$ $66.7$ $85.9$ $60.9$ $59.8$ $61.8$ $61.1$
-FAP $49.6$ $91.8$ $56.8$ $23.2$ $24.3$ $66.6$ $82.2$ $85.4$ $58.6$ $61.2$ $70.6$ $65.6$
-AdvPT $52.1$ $93.1$ $61.8$ $46.9$ $30.8$ $74.9$ $88.2$ $89.8$ $67.8$ $70.5$ $74.3$ $73.0$
-AdvPT-V $51.2$ $88.4$ $36.0$ $23.5$ $15.4$ $56.5$ $53.3$ $81.9$ $48.3$ $55.8$ $51.6$ $51.3$
-AdvPT-VLI $54.4$ $84.7$ $35.4$ $17.7$ $11.4$ $55.5$ $52.5$ $76.9$ $45.2$ $50.1$ $48.5$ $50.8$
-AdvPT-VLJ $53.0$ $86.5$ $35.2$ $16.4$ $8.5$ $52.6$ $50.7$ $77.7$ $42.5$ $51.0$ $48.6$ $49.1$
+M-Attack Vanilla $21.3$ $37.3$ $14.6$ $16.8$ $13.9$ $35.2$ $21.5$ $38.6$ $20.1$ $33.0$ $26.9$ $25.4$
+FAP $43.6$ $70.5$ $30.1$ $11.6$ $24.0$ $48.2$ $42.5$ $55.7$ $42.6$ $45.5$ $36.8$ $41.0$
+APD $41.9$ $71.6$ $25.4$ $15.7$ $19.0$ $42.3$ $36.4$ $57.0$ $41.4$ $40.3$ $33.3$ $38.6$
+AdvPT $34.4$ $43.6$ $12.8$ $8.8$ $9.6$ $38.5$ $29.0$ $47.5$ $32.6$ $39.0$ $45.8$ $31.1$
+AdvPT-V $45.3$ $67.3$ $25.7$ $13.6$ $17.8$ $46.3$ $39.5$ $56.4$ $44.4$ $39.7$ $37.4$ $39.4$
+AdvPT-VLI $47.9$ $66.4$ $31.5$ $14.8$ $16.1$ $46.7$ $40.6$ $53.2$ $49.7$ $48.0$ $45.7$ $41.9$
+AdvPT-VLJ $40.6$ $71.7$ $24.8$ $18.9$ $19.5$ $45.4$ $40.2$ $51.9$ $48.5$ $50.2$ $54.7$ $42.4$
 […]
-AGS Vanilla $53.5$ $86.8$ $30.9$ $11.0$ $18.1$ $56.9$ $56.9$ $78.0$ $54.1$ $53.7$ $49.7$ $54.3$
-FAP $51.1$ $88.2$ $50.2$ $12.3$ $22.2$ $58.9$ $77.4$ $80.1$ $53.4$ $55.0$ $61.3$ $58.5$
-AdvPT $54.6$ $88.5$ $51.5$ $26.1$ $26.1$ $55.9$ $76.0$ $80.4$ $62.6$ $62.8$ $64.1$ $61.5$
-AdvPT-V $53.0$ $85.8$ $33.2$ $15.6$ $13.6$ $47.6$ $47.7$ $78.2$ $44.3$ $51.3$ $44.3$ $50.8$
-AdvPT-VLI $56.3$ $81.1$ $33.6$ $14.7$ $10.6$ $46.2$ $48.5$ $74.3$ $41.5$ $46.2$ $41.7$ $47.5$
-AdvPT-VLJ $54.9$ $83.0$ $31.6$ $12.1$ $7.5$ $44.0$ $45.3$ $73.9$ $38.6$ $46.2$ $41.4$ $47.3$
-[…]
-PGD Vanilla $2.7$ $28.3$ $2.0$ $0.1$ $0.0$ $14.4$ $3.2$ $8.8$ $1.4$ $1.4$ $4.0$ $6.7$
-FAP $24.0$ $62.4$ $26.7$ $0.0$ $5.2$ $21.8$ $51.7$ $33.7$ $19.5$ $25.4$ $10.9$ $23.3$
-AdvPT $1.5$ $27.0$ $6.6$ $0.2$ $0.6$ $0.9$ $4.3$ $2.8$ $0.7$ $1.6$ $2.1$ $5.2$
-AdvPT-V $19.6$ $61.5$ $18.6$ $8.1$ $3.8$ $12.4$ $25.0$ $38.0$ $9.8$ $17.3$ $17.0$ $20.4$
-AdvPT-VLI $21.8$ $59.1$ $18.6$ $10.0$ $2.9$ $12.9$ $21.7$ $38.2$ $10.0$ $15.7$ $14.2$ $20.2$
-AdvPT-VLJ $20.0$ $58.7$ $16.4$ $10.2$ $2.3$ $11.4$ $20.5$ $32.8$ $8.6$ $14.9$ $12.7$ $18.9$
+PGD Vanilla $2.7$ $28.3$ $2.0$ $0.1$ $0.0$ $14.4$ $3.2$ $8.8$ $1.4$ $1.4$ $4.0$ $6.0$
+FAP $22.3$ $67.2$ $25.5$ $30.5$ $22.9$ $20.7$ $75.3$ $46.6$ $33.0$ $31.9$ $41.1$ $37.8$
+APD $23.6$ $71.4$ $31.5$ $31.0$ $23.9$ $25.2$ $74.3$ $44.9$ $48.7$ $33.3$ $40.0$ $40.7$
+AdvPT $1.5$ $27.0$ $6.6$ $0.2$ $0.6$ $0.9$ $4.3$ $2.8$ $0.7$ $1.6$ $2.1$ $4.4$
+AdvPT-V $19.6$ $61.5$ $18.6$ $8.1$ $3.8$ $12.4$ $25.0$ $38.0$ $9.8$ $17.3$ $17.0$ $21.0$
+AdvPT-VLI $21.8$ $59.1$ $18.6$ $10.0$ $2.9$ $12.9$ $21.7$ $38.2$ $10.0$ $15.7$ $14.2$ $20.5$
+AdvPT-VLJ $20.0$ $58.7$ $16.4$ $10.2$ $2.3$ $11.4$ $20.5$ $32.8$ $8.6$ $14.9$ $12.7$ $19.0$
 […]
 AutoAttack Vanilla $0.0$ $0.0$ $0.1$ $0.1$ $0.1$ $0.0$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$
-FAP $1.6$ $1.1$ $4.7$ $2.3$ $2.7$ $1.4$ $1.1$ $1.9$ $1.4$ $2.6$ $1.8$ $2.2$
+FAP $12.2$ $54.8$ $12.9$ $2.2$ $1.9$ $8.6$ $19.4$ $31.3$ $3.7$ $13.5$ $17.0$ $16.1$
+APD $10.6$ $48.2$ $14.0$ $3.6$ $2.3$ $5.8$ $13.9$ $27.5$ $4.0$ $9.9$ $10.2$ $13.6$
 AdvPT $0.1$ $0.0$ $0.2$ $0.2$ $0.4$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.0$ $0.1$
-AdvPT-V $14.9$ $55.3$ $15.0$ $2.2$ $2.2$ $8.6$ $18.8$ $32.1$ $5.9$ $12.5$ $13.0$ $17.9$
-AdvPT-VLI $17.0$ $54.3$ $14.3$ $9.2$ $1.6$ $9.1$ $16.2$ $33.6$ $4.8$ $11.4$ $11.3$ $17.0$
-AdvPT-VLJ $10.3$ $43.9$ $10.3$ $3.7$ $1.2$ $4.4$ $9.7$ $19.8$ $1.9$ $6.7$ $6.0$ $11.1$
+AdvPT-V $14.9$ $55.3$ $15.0$ $2.2$ $2.2$ $8.6$ $18.8$ $32.1$ $5.9$ $12.5$ $13.0$ $16.4$
+AdvPT-VLI $17.0$ $54.3$ $14.3$ $9.2$ $1.6$ $9.1$ $16.2$ $33.6$ $4.8$ $11.4$ $11.3$ $16.6$
+AdvPT-VLJ $10.3$ $43.9$ $10.3$ $3.7$ $1.2$ $4.4$ $9.7$ $19.8$ $1.9$ $6.7$ $6.0$ $10.7$
 […]
-TABLE II: Evaluation results on ViT-B32 under clean and adversarial settings.
-Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (CWA and AGS), where $\dagger$ presents our method.
+TABLE III: Evaluation results on ViT-B32 under clean and adversarial settings.
+Performance is reported across multiple datasets under white-box (PGD and AutoAttack) and black-box attacks (M-Attack), where $\dagger$ presents our method.
 Best results are highlighted in bold.
@@
 Aircraft
-Food101
+Food
 Flowers
@@
 […]
-CWA Vanilla $48.6$ $90.5$ $40.6$ $33.4$ $16.1$ 74.6 $61.8$ $85.2$ $55.7$ $59.5$ $60.1$ $60.1$
-FAP $50.8$ $90.4$ $55.4$ $40.1$ $21.7$ $63.9$ $81.2$ $80.6$ $48.9$ $61.5$ $62.1$ $59.7$
-AdvPT 51.2 92.9 60.0 $54.9$ $26.7$ $73.7$ $88.2$ 87.5 $65.5$ 69.3 74.3 $67.7$
-AdvPT-V $48.0$ $89.2$ $36.5$ $10.8$ $14.3$ $59.0$ $52.5$ $78.0$ $43.3$ $54.3$ $50.5$ $48.8$
-AdvPT-VLI $50.7$ $87.5$ $29.4$ $23.5$ $8.4$ $57.9$ $49.2$ $80.2$ $38.3$ $48.7$ $48.3$ $47.5$
-AdvPT-VLJ 51.2 $83.9$ $29.1$ $21.7$ $7.4$ $56.7$ $46.0$ $80.2$ $35.7$ $49.3$ $48.6$ $46.3$
 […]
-AGS Vanilla $46.1$ $85.0$ $29.8$ $11.8$ $13.4$ $50.9$ $49.1$ $73.6$ $47.5$ $49.6$ $46.9$ $45.8$
-FAP 52.0 89.9 54.4 $31.6$ $21.7$ 63.4 $81.0$ 80.1 $48.8$ 61.1 $61.5$ $58.7$
-AdvPT $48.6$ $88.0$ $48.5$ $29.9$ $23.9$ $50.5$ $71.9$ $74.9$ $57.3$ $57.5$ $60.1$ $55.6$
-AdvPT-V $47.7$ $85.9$ $33.6$ $3.6$ $12.3$ $49.3$ $46.7$ $73.6$ $37.9$ $48.9$ $44.3$ $44.0$
-AdvPT-VLI $50.6$ $83.8$ $26.4$ $14.7$ $6.7$ $49.3$ $43.0$ $76.1$ $33.9$ $43.7$ $41.8$ $42.7$
-AdvPT-VLJ $51.2$ $80.0$ $26.7$ $14.6$ $7.2$ $48.0$ $39.5$ $74.5$ $32.0$ $44.5$ $43.1$ $41.9$
-[…]
-PGD Vanilla $2.0$ $29.1$ $4.9$ $0.2$ $0.0$ $7.4$ $1.7$ $4.2$ $0.4$ $1.5$ $3.0$ $6.7$
-FAP $23.4$ $67.1$ $32.8$ $0.4$ $7.0$ $26.2$ $57.3$ $40.0$ $17.6$ $29.8$ $31.5$ $23.3$
-AdvPT $2.8$ $31.7$ $10.0$ $0.4$ $0.8$ $1.8$ $8.7$ $3.2$ $2.0$ $2.9$ $4.5$ $5.2$
-AdvPT-V $19.0$ $64.6$ $19.8$ $0.3$ $2.9$ $15.0$ $23.0$ $36.3$ $9.6$ $18.9$ $18.3$ $20.4$
-AdvPT-VLI $20.7$ $62.1$ $16.3$ $5.7$ $1.5$ $15.9$ $19.3$ $37.6$ $8.4$ $16.2$ $17.7$ $20.2$
-AdvPT-VLJ $20.9$ $59.9$ $15.2$ $10.1$ $2.0$ $15.3$ $18.8$ $37.5$ $9.6$ $16.9$ $18.2$ $18.9$
-[…]
+M-Attack Vanilla $18.3$ $33.7$ $15.0$ $11.9$ $13.1$ $31.5$ $18.7$ $35.2$ $19.5$ $28.4$ $22.6$ $22.5$
+FAP $39.7$ $65.6$ $28.3$ $9.6$ $21.5$ $48.4$ $39.8$ $51.5$ $48.8$ $40.2$ $32.0$ $38.7$
+APD $38.4$ $68.2$ $22.9$ $16.1$ $16.9$ $40.6$ $33.7$ $53.6$ $39.6$ $36.8$ $30.1$ $36.1$
+AdvPT $32.0$ $41.5$ $11.9$ $9.2$ $6.7$ $33.7$ $25.8$ $43.1$ $29.8$ $27.6$ $42.1$ $27.6$
+AdvPT-V $40.9$ $69.3$ $33.6$ $3.6$ $12.3$ $49.3$ $46.7$ $50.9$ $37.9$ $48.9$ $44.3$ $39.8$
+AdvPT-VLI $42.6$ $65.8$ $28.3$ $11.9$ $13.1$ $42.7$ $34.8$ $52.5$ $47.6$ $46.0$ $42.1$ $38.9$
+AdvPT-VLJ $42.9$ $66.1$ $28.2$ $14.8$ $9.5$ $45.6$ $33.9$ $55.8$ $48.1$ $44.7$ $45.3$ $39.5$
+NAP-Tuning 44.6 76.5 37.1 51.9 36.8 52.1 67.4 57.3 67.7 50.9 48.3 53.7
+PGD Vanilla $2.0$ $29.1$ $4.9$ $0.2$ $0.0$ $7.4$ $1.7$ $4.2$ $0.4$ $1.5$ $3.0$ $4.9$
+FAP $24.5$ $66.8$ $27.1$ $28.3$ $19.2$ $19.8$ $69.7$ $40.5$ $29.9$ $30.3$ $36.9$ $35.7$
+APD $18.4$ $66.0$ $30.8$ $22.4$ $20.2$ $26.1$ $64.5$ $33.9$ $36.8$ $30.4$ $36.8$ $35.1$
+AdvPT $2.8$ $31.7$ $10.0$ $0.4$ $0.8$ $1.8$ $8.7$ $3.2$ $2.0$ $2.9$ $4.5$ $6.3$
+AdvPT-V $19.0$ $64.6$ $19.8$ $0.3$ $2.9$ $15.0$ $23.0$ $36.3$ $9.6$ $18.9$ $18.3$ $20.7$
+AdvPT-VLI $20.7$ $62.1$ $16.3$ $5.7$ $1.5$ $15.9$ $19.3$ $37.6$ $8.4$ $16.2$ $17.7$ $20.1$
+AdvPT-VLJ $20.9$ $59.9$ $15.2$ $10.1$ $2.0$ $15.3$ $18.8$ $37.5$ $9.6$ $16.9$ $18.2$ $20.4$
+NAP-Tuning 26.5 77.9 34.2 49.1 25.4 28.4 82.4 41.3 52.4 31.5 45.0 44.9
 AutoAttack Vanilla $0.0$ $0.4$ $0.3$ $0.1$ $0.1$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.1$ $0.1$
-FAP $2.0$ $1.3$ $5.0$ $1.6$ $3.5$ $2.0$ $1.9$ $1.5$ $1.0$ $3.2$ $2.3$ $2.2$
-AdvPT $0.0$ $0.6$ $0.5$ $0.1$ $0.2$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$ $0.2$
-AdvPT-V $8.3$ $44.8$ $12.4$ $0.0$ $0.5$ $5.4$ $8.8$ $16.3$ $2.8$ $7.4$ $7.2$ $10.0$
-AdvPT-VLI $9.3$ $44.2$ $10.5$ $0.2$ $0.4$ $5.6$ $7.6$ $16.4$ $2.1$ $6.2$ $6.8$ $10.1$
-AdvPT-VLJ $10.1$ $44.7$ $10.3$ $1.2$ $0.6$ $6.1$ $7.9$ $18.9$ $2.3$ $7.2$ $8.3$ $11.0$
+FAP $10.9$ $42.8$ $8.5$ $1.8$ $1.0$ $7.1$ $16.7$ $26.2$ $2.4$ $10.0$ $12.5$ $12.7$
+APD $9.0$ $46.5$ $10.1$ $0.2$ $0.5$ $6.5$ $14.8$ $17.9$ $1.9$ $9.4$ $12.2$ $11.7$
+AdvPT $0.0$ $0.6$ $0.5$ $0.1$ $0.2$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$ $0.1$
+AdvPT-V $8.3$ $44.8$ $12.4$ $0.0$ $0.5$ $5.4$ $8.8$ $16.3$ $2.8$ $7.4$ $7.2$ $10.4$
+AdvPT-VLI $9.3$ $44.2$ $10.5$ $0.2$ $0.4$ $5.6$ $7.6$ $16.4$ $2.1$ $6.2$ $6.8$ $9.9$
+AdvPT-VLJ $10.1$ $44.7$ $10.3$ $1.2$ $0.6$ $6.1$ $7.9$ $18.9$ $2.3$ $7.2$ $8.3$ $10.7$
 […]
 V Experiments
-This section systematically evaluates the effectiveness of our proposed NAP-Tuning.
-We first introduce the experimental setup, then validate the key components of our method through a series of carefully designed experiments, and comprehensively compare with existing approaches.
 V-A Experimental Setup
@@
 For the ImageNet test set, consistent with prior adversarial attack studies [4, 29, 33], we randomly sample 1,000 images, ensuring one image per class.
-Our experiments are conducted on the CLIP model, where the default configuration employs the publicly available ViT-B/16 architecture unless otherwise specified.
-We also include results for the ViT-B/32 model as a supplementary evaluation.
-In accordance with previous work, we use hand-crafted prompts as textual inputs (e.g., “a photo of a <class>, a type of pet” for Oxford Pets).
-V-A2 Attack Methods
-To evaluate adversarial robustness, we implement both white-box and black-box adversarial attacks with more stringent settings compared to our conference version.
-For white-box attacks, we use learnable textual vectors to construct cross-entropy loss and employ stronger PGD-100 [4] and AutoAttack [35].
-For black-box attacks, we utilize more advanced methods such as CWA [36] and AGS [37].
-V-A3 Defense Baselines
-We compare our method against several strong baselines, including variants of AdvPT and the recent FAP [5]:
-  • AdvPT: the original Adversarial Prompt Tuning method, which performs prompt tuning exclusively on the text side.
-It adapts the textual input prompts to improve robustness without modifying the visual inputs.
-  • AdvPT-V: an extension of AdvPT that applies prompt tuning on the vision side.
-Specifically, it introduces additional learnable prompts into the visual encoder, enabling adaptation of the visual input space to adversarial perturbations.
-  • AdvPT-VLJ (vision-language joint prompt tuning): this variant simultaneously introduces learnable prompts to both the vision and text modalities.
-Moreover, it establishes a mapping between the vision and language prompts, allowing joint adversarial tuning of both modalities.
-This structure follows a design similar to MaPLe [8].
-  • AdvPT-VLI (vision-language independent prompt tuning): a modification of AdvPT-VLJ that removes the mapping between vision and language prompts.
-In this setup, the learnable prompts for the visual and textual inputs are optimized independently, without explicit cross-modal interactions.
-  • FAP [5]: a recent method proposed for few-shot adversarial robustness, which carefully designs the objective function.
-V-A4 Implementation Details
-Our training framework spans 90 epochs with a batch size of 512.
+Our experiments are conducted on the CLIP model.
+In accordance with previous work, we use hand-crafted prompts as textual inputs (e.g., “a photo of a <class>, a type of pet” for Pets).
+To evaluate adversarial robustness, we implement both white-box and black-box adversarial attacks.
+For white-box attacks, we use PGD-100 [16] and AutoAttack [3].
+Consistent with our training protocol, we use a perturbation magnitude of $\epsilon=1/255$.
+This provides a direct evaluation of the robustness learned at the training budget.
+For black-box attacks, we utilize the strong M-Attack [11].
+As black-box attacks are inherently less effective, and to ensure a challenging and meaningful evaluation, we follow standard evaluation practice and use a larger perturbation budget of $\epsilon=8/255$.
+V-A2 Defense Baselines
+We compare NAP-Tuning against state-of-the-art prompt-based defenses.
+These include the original AdvPT (text-only) and its multi-modal extensions: AdvPT-V (vision-only), AdvPT-VLI (independent dual-modal), and AdvPT-VLJ (joint mapping akin to MaPLe [8]).
+We also include recent strong baselines: FAP [41], which optimizes a few-shot robustness objective, and APD [15], which utilizes bimodal distillation from a clean teacher.
+V-A3 Implementation Details
+Our training framework spans 90 epochs with a batch size of 32.
 We adopt the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is modulated through a cosine annealing schedule.
-For adversarial example generation during training (as defined in Equation (12)), we implement PGD-5 with a default perturbation magnitude $\epsilon=1/255$ for both training and testing phases, though we analyze the effects of different $\epsilon$ values in subsequent sections.
+For adversarial example generation during training, we implement PGD-5 with a default perturbation magnitude $\epsilon=1/255$ for training phases, though we analyze the effects of different $\epsilon$ values in subsequent sections.
 For the TokenRefiner $R$, we implement a default 2-layer neural network architecture, with the impact of network depth examined in our ablation studies.
@@
 The hyperparameter $\alpha_{0}$ in Equation (17) is set to 5.0.
-All experiments are conducted on NVIDIA A800 80GB GPUs.
-V-B Main Results
-The main evaluation results are summarized in Tables I and II.
+All experiments are conducted on a single NVIDIA A800 80GB GPU.
+V-B Main Results and Comparisons
+We begin by evaluating the overall performance of NAP-Tuning against existing methods.
+We first compare it to state-of-the-art prompt-based defenses to establish its effectiveness, and then benchmark it against fine-tuning based approaches to highlight the architectural trade-offs.
+V-B1 Comparison with State-of-the-Art Baselines
+The main evaluation results are summarized in Tables II and III.
 Overall, NAP-Tuning achieves highly competitive clean accuracy compared to existing baselines.
 While its clean performance is slightly lower than that of AdvPT, this is expected, as AdvPT explicitly prioritizes clean accuracy in the robustness-accuracy trade-off, often at the cost of adversarial robustness.
-In terms of adversarial robustness, NAP-Tuning demonstrates clear and consistent advantages.
-Under black-box attacks, particularly CWA and AGS, our method outperforms all baselines across both ViT-B16 and ViT-B32 backbones, highlighting its superior generalization to unseen adversarial threats.
-More importantly, under white-box attacks such as PGD and AutoAttack, NAP-Tuning exhibits significant improvements.
-AutoAttack is widely regarded as the most rigorous and comprehensive white-box evaluation benchmark, combining several strong attacks to provide a thorough assessment of model robustness.
-Under this challenging setup, NAP-Tuning achieves substantial gains over all competitors: on ViT-B16, it attains 51.4% on average under AutoAttack, outperforming the strongest baseline (AdvPT-V) by 33.5%; on ViT-B32, it reaches 44.0%, exceeding the best competitor (AdvPT-VLJ) by 33.0%.
-These results clearly demonstrate the robustness of NAP-Tuning against powerful, diverse adversarial threats.
-V-C Multi-layer Prompt Architecture
+NAP-Tuning demonstrates consistent advantages across scenarios.
+Under the black-box M-Attack, it shows superior generalization, surpassing the strongest baselines by 16.4% on ViT-B16 and 13.9% on ViT-B32.
+This lead is further amplified under strenuous white-box attacks (e.g., PGD, AutoAttack).
+Notably, on the rigorous AutoAttack benchmark, our method achieves 48.9% on ViT-B16 and 44.0% on ViT-B32 (see Tables II and III), significantly outperforming all competitors.
+V-B2 Comparison with Fine-tuning Defense Methods
+We compare NAP-Tuning with TeCoA [18], which fine-tunes CLIP’s encoder layers for adversarial robustness.
+This comparison directly addresses the fundamental question: should we modify the pre-trained encoder or keep it frozen?
+Both methods are evaluated on ViT-B32 with adversarial examples generated using PGD-100.
+Figure 4: Performance comparison between NAP-Tuning (frozen image encoder) and TeCoA (fine-tuned image encoder).
+Impact of Fine-tuning.
+Fig.
+4 reveals a critical trade-off.
+While TeCoA achieves a smaller robustness gap (15.3% vs.
+24.2%), it suffers a catastrophic 32.3% drop in clean accuracy (69.1% $\to$ 36.8%) compared to NAP-Tuning.
+This failure, consistent across datasets like EuroSAT, confirms that fine-tuning the backbone disrupts the pre-trained multi-modal alignment essential for zero-shot capabilities.
+In contrast, NAP-Tuning preserves the encoder’s generalization ability by using the Neural Augmentor for targeted feature purification, thereby maintaining strong performance on both clean and adversarial data without compromising the pre-trained representations.
+TABLE IV: Training time comparison on Pets dataset. All methods trained for the same number of epochs with identical batch size on a single NVIDIA A800 GPU.
+Method Training Efficiency
+Time (minutes) Relative
+Input-level Prompt Methods
+[…]
+Deep Prompt Methods
+[…]
+[…]
+[…]
+[…]
+[…]
+Backbone Fine-tuning Methods
+[…]
+[…]
+V-C Component Ablation Studies
+Having demonstrated NAP-Tuning’s superior performance, we now conduct a series of ablation studies to deconstruct our framework.
+We systematically evaluate the impact of each core design choice, from the high-level prompt architecture down to the specific configuration of our purification module.
+V-C1 Effect of Multi-layer Prompt Architecture
+Figure 5: Clean and robust accuracy across datasets when varying the number of prompt layers (1-12).
+Complex datasets (ImageNet, Food101, SUN397) show optimal performance at intermediate depth, while other datasets benefit from deeper architectures.
 To investigate how the depth of our multi-layer prompt architecture affects model robustness, we conduct an ablation study by varying the number of layers where prompts are inserted, from 1 to 12.
@@
 5 presents the performance on eleven diverse datasets under both clean and adversarial conditions.
-The results reveal several key insights.
-First, as shown in the average performance (Fig.
-4l), both clean and robust accuracy generally improve as the number of layers increases, confirming the effectiveness of our hierarchical prompting strategy.
-Interestingly, we observe dataset-specific patterns related to task complexity.
-For more challenging datasets such as ImageNet (Fig.
-4f), SUN397 (Fig.
-4j), and Food101 (Fig.
-4e), performance follows an inverted U-shape, peaking at intermediate layer depths before declining.
-This pattern indicates mild overfitting when increasing parameter capacity without corresponding increases in training examples (shot=16).
-The highlighted regions in these plots mark the optimal layer ranges for these datasets.
-These findings reveal the potential for further performance enhancements through larger training sets and dataset-specific layer optimization, suggesting that improved robustness can be achieved through straightforward data augmentation.
-Nevertheless, our method demonstrates strong performance even with the default configuration (layer=12) as shown in Table I and Table II, consistently outperforming baseline approaches across diverse datasets.
-Figure 4: Clean and robust accuracy across datasets when varying the number of prompt layers (1-12).
-Complex datasets (ImageNet, Food101, SUN397) show optimal performance at intermediate depth, while other datasets benefit from deeper architectures.
-V-D Adversarial Regularization Parameter
-We analyze the impact of adversarial regularization weight ($\alpha$) on both standard (clean) and adversarial (robust) accuracy.
 Fig.
-LABEL:fig:alpha presents results across $\alpha$ values from 0.01 to 100, with panel (a) showing both metrics on a single y-axis and panel (b) using dual y-axes to highlight relative trends.
-The results show that clean accuracy remains stable (98.5%-99.1%) across all $\alpha$ values, while robust accuracy increases substantially from 79.9% at $\alpha$=0.01 to 93.9% at $\alpha$=2.0, with diminishing gains thereafter.
-For $\alpha\geq 5.0$, robust accuracy consistently exceeds 94%.
-Rather than specifically optimizing for the peak performance at $\alpha$=20.0 (94.4% robust accuracy), we adopt $\alpha$=5.0 as our representative configuration, which achieves 98.8% clean accuracy and 94.2% robust accuracy.
-These findings demonstrate that our approach effectively balances clean accuracy and adversarial robustness across a broad range of $\alpha$ values, exhibiting strong performance without requiring fine-tuned parameter selection.
-V-E TokenRefiner Architecture
-We analyze the impact of TokenRefiner network depth on model performance using Oxford Flowers (simpler task) and ImageNet (more challenging task).
-Fig.
-LABEL:fig:token presents clean and robust accuracy across different TokenRefiner layer configurations.
-Results demonstrate that TokenRefiner is essential for model convergence—its absence (layer = 0) prevents the model from converging under 5-step PGD training, resulting in poor performance (4.6% clean accuracy on ImageNet, 25.8% on Oxford Flowers).
-Performance improves substantially as layers increase from 0 to 2, with ImageNet clean accuracy rising from 4.6% to 54.4% and robust accuracy from 1.9% to 28.0%.
-Similarly, Oxford Flowers shows improvements from 25.8% to 96.1% (clean) and from 15.1% to 86.6% (robust).
-Performance stabilizes beyond 2 layers, leading us to select a 2-layer architecture for our final implementation.
-Notably, unlike previous approaches such as FAP [5] and APD [6], which lack specialized prompt structure designs and are therefore limited to weaker adversarial search strategies, our TokenRefiner enables convergence under stronger adversarial perturbations.
-The consistent pattern across datasets of varying complexity suggests our approach generalizes well without requiring task-specific architectural adjustments.
-V-F Perturbation Magnitude
-We analyze the impact of perturbation budgets during both training and testing phases.
-Table III shows results on the Flowers dataset with varying $\epsilon$ values.
-The results reveal a clear trade-off between clean accuracy and adversarial robustness.
-As training perturbation budget increases from $\epsilon=1$ to $\epsilon=4$, clean accuracy decreases from 96.1% to 87.5%, while robustness against stronger attacks improves substantially.
-Models trained with larger $\epsilon$ values perform significantly better when tested under matching or higher perturbation budgets—a model trained with $\epsilon=4$ achieves 55.9% accuracy when tested at the same level, versus only 17.1% for a model trained with $\epsilon=1$.
-Models experience rapid performance degradation when tested with perturbations larger than their training budget.
-This observation suggests that, similar to traditional adversarial training, adversarial prompt tuning exhibits a clear trade-off between clean accuracy and adversarial robustness.
-TABLE III: Accuracy (%) under $\ell_{\infty}$-bounded adversarial perturbations.
-Models are trained and evaluated with varying perturbation magnitudes (measured in $\epsilon/255$).
-Bold values indicate the highest robustness at each test-time perturbation level.
+5 confirms that increasing prompt depth generally enhances both clean and robust accuracy.
+However, complex datasets (e.g., ImageNet, Food101) exhibit an inverted U-shape trajectory, peaking at intermediate depths.
+This pattern likely stems from mild overfitting given the limited training samples (16-shot).
+Despite these task-specific variations, the full-depth configuration (layer=12) yields the superior average performance and consistently outperforms baselines, justifying its adoption as the default setting.
+V-C2 Effect of the Neural Augmentor Module
+Figure 6: Ablation analysis comparing models with and without the Neural Augmentor (NA) across varying prompt depths. The complete model consistently outperforms the ablated variant.
+Complementing the multi-layer analysis, Fig.
+6 isolates the specific contribution of the Neural Augmentor.
+The results unequivocally show that the complete model yields consistently higher robust accuracy across all layer configurations on both ImageNet and Oxford Flowers.
+A particularly interesting observation is the interaction between module capacity and task complexity.
+On ImageNet, the complete model peaks earlier (around Layer 4) compared to the ablated variant, which plateaus much later (around Layer 9).
+This shift suggests that the Neural Augmentor accelerates adaptation by providing necessary structural capacity.
+While this added capacity introduces a mild overfitting tendency in few-shot settings (manifesting as the inverted U-shape), the substantial performance gain confirms that the benefits of explicit feature purification far outweigh these limitations.
+V-C3 Effect of TokenRefiner Modality
+Figure 7: Modality-specific TokenRefiner ablation study under PGD-100 attacks.
+While image-side TokenRefiner provides substantial robustness benefits, the dual-modal approach consistently outperforms single-modality configurations across all datasets.
+We analyze the modality-specific contributions in Fig.
+7.
+A clear hierarchy emerges: the Image-only TokenRefiner serves as the primary defense backbone, driving the vast majority of robustness gains.
+This aligns with the theoretical expectation that direct visual purification is essential against image-based attacks.
+Interestingly, the Text-only variant—despite operating on unperturbed inputs—still yields considerable robustness improvements.
+This suggests a mechanism of cooperative feature alignment, where the module dynamically adapts clean text embeddings to better match the purified visual features.
+Consequently, the full Dual-modal framework consistently outperforms single-branch configurations, confirming that text-side alignment is a critical, non-redundant complement to visual purification.
+V-C4 Effect of TokenRefiner Architecture Depth
+We investigate the impact of TokenRefiner network depth on model performance using Oxford Flowers (simpler task) and ImageNet (complex task) in Fig.
+9(a).
+The results reveal that the TokenRefiner is not merely an enhancement but a prerequisite for optimization stability.
+Specifically, removing the module (layer = 0) leads to model collapse under 5-step PGD adversarial training, resulting in near-random performance on both datasets.
+However, increasing the depth to just 2 layers resolves this instability, yielding dramatic improvements that restore both clean and robust accuracy to competitive levels.
+Beyond this point, performance saturates, justifying our selection of a 2-layer architecture as the optimal configuration.
+Notably, this structural stability distinguishes our approach from previous methods like FAP [41] and APD [15]; whereas those methods suffer from convergence issues under strong adversarial search, our TokenRefiner enables robust optimization even under rigorous perturbation constraints.
+This consistent behavior across datasets of varying complexity confirms that our design generalizes well without necessitating task-specific tuning.
+(a) TokenRefiner performance across different network depths. Performance improves dramatically from 0 to 2 layers before stabilizing, with consistent patterns across datasets of varying complexity.
+V-D In-depth Analysis
+Beyond ablating components, we now provide a deeper analysis of our method’s core mechanism.
+We investigate how NAP-Tuning works by visualizing its effect on feature space, and explore its compatibility with other defense paradigms.
+V-D1 Analysis of Feature Purification
+To investigate the effectiveness of different defensive strategies in preserving feature integrity under adversarial attack, we analyze both quantitative distortion measurements and qualitative attention visualizations.
+TABLE V: Feature distortion magnitude across different defensive approaches measured using Equation (9). Lower value is better.
+Method Feature Distortion ($L_{2}$ Norm)
+Vanilla 8.67
+AdvPT 8.63
+APD 5.01
+AdvPT-V 4.87
+AdvPT-VLI 4.76
+FAP 3.58
+NAP-Tuning 1.96
+TABLE VI: Accuracy (%) under $\ell_{\infty}$-bounded adversarial perturbations. Models are trained and evaluated with varying perturbation magnitudes (measured in $\epsilon/255$).
 Training $\bm{\epsilon}$ Clean Accuracy Robust Accuracy @ Test-time $\bm{\epsilon}$
@@
 4 87.5 79.2 75.7 55.9 11.8
-V-G Few-shot Learning
+Refer to caption Refer to caption Refer to caption Refer to caption
+Original Adversarial AdvPT AdvPT-V
+Refer to caption Refer to caption Refer to caption Refer to caption
+AdvPT-VLI APD FAP NAP-Tuning
+Figure 9: EigenCAM visualization of attention patterns.
+NAP-Tuning maintains focus on semantically meaningful regions (cat’s face) even under adversarial attack, while other methods exhibit significant attention shifts toward irrelevant areas.
+Table V quantifies the feature distortion magnitude under attack.
+A distinct capability hierarchy emerges from the results.
+Shallow approaches (Vanilla CLIP, AdvPT) exhibit significant distortion, confirming that input-level optimization alone fails to restrain feature corruption.
+Methods incorporating deeper interventions (e.g., APD, AdvPT-V) achieve moderate mitigation.
+In sharp contrast, NAP-Tuning outperforms all competitors by a wide margin, achieving the lowest distortion score.
+This represents a drastic reduction compared to the baseline, validating our method’s unique ability to maintain structural integrity in the embedding space.
+These quantitative findings are strongly corroborated by the EigenCAM [19] visualizations in Fig.
+9.
+While baseline models suffer from severe attention dispersion—scattering focus across irrelevant background regions—NAP-Tuning maintains precise semantic consistency.
+Specifically, our model retains focus on discriminative regions (e.g., facial features) even under adversarial attack, closely mirroring the attention patterns of clean inputs.
+This alignment between reduced feature distortion and preserved attention focus confirms our central hypothesis: adversarial vulnerability stems primarily from internal feature corruption, which the Neural Augmentor effectively rectifies.
+V-D2 Analysis of Model-Image Defense Integration
+We investigate NAP-Tuning’s capability to integrate with image-based defense mechanisms.
+Similar to AdvPT [37], our approach focuses exclusively on model-side adaptation, allowing it to complement image-based defense techniques.
+This presents an advantage over methods that already incorporate image transformations in their design, which often cannot be effectively combined with other preprocessing defenses.
+Figure 10: Performance comparison across eleven datasets demonstrates consistent improvements when combining model-centric NAP-Tuning with the image-processing DiffPure method.
+Figure 11: Effect of adversarial regularization weight ($\alpha$) on clean and robust accuracy on Flowers dataset.
+To demonstrate this complementary capability, we evaluate NAP-Tuning combined with DiffPure [20], a diffusion-based adversarial purification technique.
+Fig.
+10 shows the combined approach consistently outperforms standalone NAP-Tuning, with an average improvement of 12.9 percentage points.
+These results highlight a key advantage of our model-focused approach: NAP-Tuning enhances network parameter robustness while remaining fully compatible with input-processing techniques.
+This orthogonality enables effective integration that addresses different aspects of adversarial vulnerability.
+The consistent improvements across diverse datasets suggest that combining complementary defense strategies represents a promising direction for building more robust visual recognition systems.
+V-E Parameter Sensitivity Analysis
+Finally, we evaluate the sensitivity of NAP-Tuning to key hyperparameters and training settings.
+These experiments validate the robustness and stability of our approach across different configurations.
+V-E1 Impact of Adversarial Regularization ($\alpha$)
+We investigate the sensitivity of NAP-Tuning to the adversarial regularization weight ($\alpha$) in Fig.
+11.
+The analysis reveals a highly favorable decoupling of metrics: clean accuracy remains remarkably stable across the entire parameter sweep, showing negligible fluctuation regardless of the regularization strength.
+In contrast, robust accuracy exhibits a rapid ascent as $\alpha$ increases from minimal values, effectively stabilizing around $\alpha=2.0$.
+Beyond this threshold ($\alpha\geq 5.0$), performance saturates at a high level with diminishing returns.
+Consequently, we adopt $\alpha=5.0$ as our default configuration.
+This setting yields a near-optimal balance without strictly necessitating the absolute peak observed at higher values (e.g., $\alpha=20.0$).
+These findings confirm that our approach is insensitive to precise hyperparameter tuning, maintaining consistent effectiveness across a broad operational range.
+Figure 12: Impact of shot count on clean and robust accuracy across eleven datasets.
+Each subplot (a-k) shows performance on individual datasets, while subplot (l) presents the average across all datasets.
+Figure 13: Effect of context vector count (1-16) on ImageNet and Oxford Flowers.
+Performance varies minimally across different vector counts, indicating context vectors are not a primary bottleneck for model capability.
+V-E2 Impact of Perturbation Magnitude ($\epsilon$)
+We evaluate the interplay between perturbation budgets and model performance in Table VI.
+A clear trade-off emerges from the data: increasing the training perturbation magnitude ($\epsilon$) substantially bolsters robustness against stronger attacks, yet this hardening incurs a discernible cost in clean accuracy.
+Specifically, models trained with larger budgets demonstrate vastly superior resilience when subjected to matching or higher-intensity threats.
+In contrast, models trained on weaker perturbations suffer from rapid degradation when facing attacks that exceed their training budget.
+This aligns with well-established observations in traditional adversarial training, confirming that robust prompt tuning follows the same fundamental principle: resistance to extreme perturbations is achieved by sacrificing a degree of standard generalization.
+V-E3 Impact of Few-shot Count
 We examine the impact of shot count on model performance.
@@
 12 illustrates the relationship between number of shots (1-16) and model accuracy across diverse datasets.
-Clean accuracy shows consistent improvement, with average accuracy increasing from 40.6% (single shot) to 73.2% (16 shots)—a 1.8× improvement.
-More remarkably, robust accuracy exhibits steeper relative gains, rising from 14.1% to 49.8%—a dramatic 3.5× improvement.
+Clean accuracy shows consistent improvement, with average accuracy increasing from 40.6% (single shot) to 73.5% (16 shots).
+More remarkably, robust accuracy exhibits steeper relative gains, rising from 14.1% to 49.9%.
 This differential impact reveals a key insight: while few-shot learning is known to enhance clean accuracy, its benefits for robust performance are substantially more pronounced.
 The results indicate that robustness requires a richer representation of class concepts that becomes increasingly available with additional examples.
-These findings align with our observations in Section V-C, confirming that increasing shot count offers tremendous potential for performance enhancement in our method, providing a straightforward path to significantly improved robustness without architectural modifications.
-Figure 6: Impact of shot count on clean and robust accuracy across eleven datasets.
-Each subplot (a-k) shows performance on individual datasets, while subplot (l) presents the average across all datasets.
-V-H Context Vector Count
-We examine how the number of learnable context vectors ($\mathcal{V}_{t}$) affects model performance.
-Fig.
-7 shows results for models with 1-16 context vectors on ImageNet and Oxford Flowers datasets.
-For ImageNet, clean accuracy remains relatively stable (55.9%-58.1%) across different vector counts, while robust accuracy shows modest improvements with increased vectors, peaking at 29.6% with 14-16 vectors.
-On Oxford Flowers, both metrics maintain high performance across all configurations, with clean accuracy ranging from 95.5% to 96.5% and robust accuracy between 85.2% and 86.7%.
-These findings suggest that the number of context vectors $\mathcal{V}_{t}$ has a relatively minor impact on overall performance compared to the influence of layer count in Multi-Layer Prompts observed in Section V-C.
-This indicates that input-level representational capacity is not the primary performance bottleneck, particularly for less complex datasets where even few context vectors prove sufficient.
-Figure 7: Effect of context vector count (1-16) on ImageNet and Oxford Flowers.
-Performance varies minimally across different vector counts, indicating context vectors are not a primary bottleneck for model capability.
+These findings align with our observations in Section V-C1, confirming that increasing shot count offers tremendous potential for performance enhancement in our method, providing a straightforward path to significantly improved robustness without architectural modifications.
+V-E4 Impact of Context Vector Count
+We investigate the sensitivity of model performance to the length of learnable context vectors ($\mathcal{V}_{t}$) in Fig.
+13.
+On the complex ImageNet dataset, clean accuracy remains remarkably stable regardless of vector length, while robust accuracy exhibits only marginal gains as the number of vectors increases, eventually saturating at higher counts.
+This stability is even more pronounced on the simpler Flowers dataset, where both metrics maintain consistently high performance with negligible variance across all configurations.
+Crucially, these fluctuations are minimal when compared to the significant impact of prompt depth observed in Section V-C1.
+This confirms that input-level representational capacity is not the primary bottleneck; rather, the structural depth of the intervention plays a far more decisive role in determining robustness than the mere quantity of input tokens.
 VI Conclusion and Discussion
-This paper presents the Neural Augmentor framework for Multi-modal Adversarial Prompt Tuning (NAP-Tuning), a novel architecture-focused approach to enhancing adversarial robustness in Vision-Language Models.
-Our work addresses significant limitations in existing adversarial prompt tuning methods by fundamentally reconceptualizing prompt architecture through feature purification mechanisms.
-VI-A Conclusion
-We have extended adversarial prompt tuning from text-only to multi-modal contexts and from single-layer to multi-layer architectures while introducing token refiners that enable modality-specific and layer-specific feature correction.
-This architectural innovation represents a paradigm shift from previous loss-focused approaches, emphasizing the critical role of feature purification in addressing adversarial distortions.
-Through extensive experimentation, we have demonstrated that NAP-Tuning substantially outperforms existing methods, achieving state-of-the-art robustness against strong attacks while maintaining competitive clean accuracy.
-VI-B Broader Implications
-Our findings have several significant implications: 1) highlight the insufficiency of merely transplanting prompt tuning techniques from generalization domains to adversarial defense without architectural reconsideration; 2) underscore the importance of addressing adversarial perturbations at the feature level rather than solely through loss modification.
-VI-C Discussion
-Future work could explore adaptive token refiners that dynamically adjust their behavior based on detected perturbation characteristics, integration of uncertainty quantification mechanisms to improve refinement reliability, and extension to other multi-modal tasks beyond classification.
-Additionally, investigating the theoretical connections between feature purification and information bottleneck principles could yield deeper insights into adversarial robustness in representation learning.
-In conclusion, our work establishes that addressing adversarial robustness in VLMs requires fundamental architectural innovations rather than merely adapting existing prompt tuning techniques.
-The Neural Augmentor framework provides a principled approach to this challenge, offering both improved performance and conceptual insights that advance our understanding of robust multi-modal learning.
+This paper presents the Neural Augmentor framework (NAP-Tuning) as a modular structural augmentation for enhancing the adversarial robustness of Vision-Language Models.
+By introducing internal TokenRefiners into a coordinated multi-modal and multi-layer prompting structure, we enable the explicit rectification of adversarial distortions within the feature manifold.
+Our empirical results across 11 benchmarks demonstrate that this internal purification mechanism significantly outperforms conventional input-side alignment methods, particularly under rigorous evaluation protocols.
+Crucially, the plug-and-play nature of the Neural Augmentor ensures high compatibility with frozen VLM backbones, providing a scalable and parameter-efficient defense for robust multi-modal learning.
+Our analysis underscores that adversarial vulnerability in VLMs stems primarily from internal feature distortions that input-side prompt tuning alone cannot fully rectify.
+By shifting the defense focus to structural feature-level intervention, NAP-Tuning provides a principled way to restore representational integrity.
+Future research will focus on developing adaptive TokenRefiners capable of dynamic adjustment based on perturbation intensity, as well as extending this purification logic to generative multimodal architectures beyond contrastive learning paradigms.